### v1

# Mixed Hodge structures for vanishing cycles and orbifold cohomology

Abstract
Above a Laurent polynomial $f$ one makes grow a vector space of vanishing cycles (after the work of Sabbah, singularity setting), a graded Milnor ring (after the work of Kouchnirenko) and an orbifold cohomology ring (after the work of Borisov, Chen and Smith). Under suitable assumptions, these structures are isomorphic and these identifications are interesting because some results are more explicit in one setting than in another. In particular, and in order to understand better the real structures and the dualities appearing in the singularity setting, we first look for the counterpart of Sabbah’s mixed Hodge structures, initially defined on the space of vanishing cycles, on the orbifold cohomology ring. Then, we discuss to what extent the orbifold Poincaré duality defined by Chen and Ruan provides a polarization of this mixed Hodge structure. We study in details the Hodge-Tate case, which can be read off from the ages of the sectors, a variation of the hard Lefschetz condition introduced by Fernandez. These notes go along with prior works of Fernandez and Wang.
1 Introduction
Let $f$ be a convenient and nondegenerate Laurent polynomial on $(\mathbb{C}^{*})^{n}$ in the sense of Kouchnirenko [22]. One associates with $f$:
  • a mixed Hodge structure $\operatorname{MHS}_{f}=(H,F^{\bullet},W_{\bullet})$ (that we call Sabbah’s mixed Hodge structure, see [28], [29]) where $H$ denotes the space of vanishing cycles; the weight filtration $W_{\bullet}$ is the weight filtration of a nilpotent endomorphism of $H$ (the logarithm of the unipotent part of the monodromy at infinity of $f$) and the Hodge filtration $F^{\bullet}$ has a very simple description via the Brieskorn lattice of $f$ and a suitable Kashiwara-Malgrange $V$-filtration,
  • the $\mathbb{Q}$-graded Milnor ring $A^{*}:=\operatorname{gr}^{\mathcal{N}}_{*}(\mathbb{C}[u,u^{-1}]/(u\partial f/\partial u))$, where $\mathcal{N}_{\bullet}$ denotes a Newton filtration suitably normalized and $(u\partial f/\partial u)$ denotes the ideal generated by the partial derivatives $(u_{1}\partial f/\partial u_{1},\ldots,u_{n}\partial f/\partial u_{n})$ of $f$ (the basic reference for this framework is Kouchnirenko’s paper [22]),
  • if the Newton polytope $P$ of $f$ is simplicial, the orbifold cohomology ring $H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P}):=H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P},\mathbb{C})$ equipped with the orbifold cup-product, of a suitable toric Deligne-Mumford stack $\mathcal{X}_{P}$ in the sense of [2] (by loc. cit., a toric Deligne-Mumford stack corresponds to a stacky fan, that is a simplicial fan with a distinguished lattice point on each ray of the fan; in order to define $\mathcal{X}_{P}$, we consider the stacky fan whose simplicial fan $\Sigma_{P}$ is the one over the faces of $P$ and whose lattice points on the rays are the vertices of the polytope $P$).
Sabbah’s mixed Hodge structures have some interesting applications, which justify the interest shown in it: see for instance [8], [12], [13], [14], [24] for the relations with Frobenius manifolds and mirror symmetry. By [28, Theorem 4.3], it is also isomorphic to the limit mixed Hodge structure constructed by Steenbrink and Zucker in [39] on $H^{\bullet}(U,f^{-1}(t),\mathbb{Q})$ when $t\rightarrow\infty$.
The first part of this text surveys the relationships between these three settings. On the one hand, if
$f(u)=\sum_{b\in\mathcal{V}(P)}u^{b}$
where $\mathcal{V}(P)$ denotes the set of the vertices of the simplicial polytope $P$ (in what follows, we always assume that $f$ has this form), we have an isomorphism of graded rings
$A^{*}\cong H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P})$ (1)
where $H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P})$ is equipped with the orbifold cup-product, see [2, Theorem 1.1]. On the other hand, and thanks to the definition of the Hodge filtration of the mixed Hodge structure $\operatorname{MHS}_{f}$, we have an isomorphism
$\operatorname{gr}_{F}^{p}H\cong\bigoplus_{n-1-p<\alpha\leq n-p}A^{\alpha}$
for $p\in\mathbb{Z}$. Using the characteristic properties of the mixed Hodge structures, we finally get isomorphisms
$H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}A^{*}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P}),$ (2)
see Section 2.5. These isomorphisms are useful because some structures are described more easily in one setting than in another: the ring structure is obviously simpler on $A^{*}$, while the linear structure is more explicit on $H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P})$, thanks to the descriptions given in [2] with the help of Stanley-Reisner rings. The first motivation of these notes is to use these isomorphisms in order to understand better the real structure and the dualities appearing in the singularity theory setting (basically provided by the Lefschetz thimbles [27] and K. Saito’s higher residue pairings [31]), by constructing explicit analogues in the orbifold cohomology setting.
We first look for the image of the mixed Hodge structure $\operatorname{MHS}_{f}$ in $H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P})$. Using the direct sum decomposition
$H_{\operatorname{orb}}^{2\alpha}(\mathcal{X}_{P}):=\bigoplus_{\ell\in F}H^{2(\alpha-i_{\ell})}(X_{\ell})$ (3)
where $i_{\ell}$ denotes the age of the sector $X_{\ell}$ and $F$ labels the set of the sectors, we define in Section 3.1 a Hodge filtration and a weight filtrations on the orbifold cohomology, which correspond to the filtrations of the mixed Hodge structure $\operatorname{MHS}_{f}$ via the isomorphisms (2). Originally, these filtrations where inspired by the description given in [14] of the mixed Hodge structure $\operatorname{MHS}_{f}$ for mirror partners of weighted projective spaces and match with the ones defined in a different setting by Wang [42].
The last ingredient that we need in order to define a mixed Hodge structure is a conjugation on $H_{\operatorname{orb}}^{2*}(\mathcal{X})$. Quite remarkably, it turns out that, once the Hodge filtration and the weight filtration are defined, the correct ones to consider on $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ are those which satisfy
$\overline{H^{2i}(X_{\ell})}\subset\oplus_{j\geq i}H^{2j}(X_{\ell^{-1}})$
for each sector $X_{\ell}$, where $\ell^{-1}$ denotes the image of $\ell$ under a natural involution on the labels (see Section 3.2). We will call such conjugations Poincaré-Lefschetz conjugations. We show in Theorem 3.10 that each Poincaré-Lefschetz conjugation gives rise to a mixed Hodge structure on $H_{\operatorname{orb}}^{2*}(\mathcal{X})$: we will denote by $\operatorname{MHS}_{\operatorname{orb}}$ such a mixed Hodge structure.
Then, we compute the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{\operatorname{orb}}$ and $\operatorname{MHS}_{f}$ in terms of the Betti numbers of the sectors $X_{\ell}$, see Proposition 3.12. Let us point out that, keeping in mind the identification between Sabbah’s mixed Hodge structure and the limit mixed Hodge structure on $H^{\bullet}(U,f^{-1}(t),\mathbb{Q})$ alluded to above, the computation of such Hodge numbers is also considered in other papers using different techniques (basically, motivic Milnor fibers instead of Brieskorn lattice), see for instance [20], [25] and the survey [40].
It follows immediately that the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ is of Hodge-Tate type if and only if
[…]
(an analogous statement is shown in [9], using a different method). Naturally enough, we will call (4) the Hodge-Tate condition. When the orbifold cohomology ring is integrally graded, this condition coincides with the one considered by Fernandez in [15, (4.2)]. We analyze some of its consequences in Section 4. If the orbifold cohomology is integrally graded and if (4) is satisfied, we have $\operatorname{gr}^{W}_{2k}H^{2*}_{\operatorname{orb}}(\mathcal{X})=H^{2(n-k)}_{\operatorname{orb}}(\mathcal{X})$ where $W_{\bullet}$ is the weight filtration of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$: in this way, we get Hodge structures on $H^{2(n-k)}_{\operatorname{orb}}(\mathcal{X})$ which we compare with the one considered in [15].
In a different combinatorial framework, a similar study is carried out by Wang in [42] using the mixed Hodge structure defined on the ring of conewise linear functions by McMullen [26] (see also [16]). Our filtrations are analogous to the ones considered in loc. cit. The corresponding formulas for the Hodge numbers, although stated in different mathematical settings (Wang’s formula is expressed in terms of the $h$-vector of the polytope $P$), complement each other: depending on the context, one formula could be applied rather than another. However, it should be emphasized that the mixed Hodge structures $\operatorname{MHS}_{\operatorname{orb}}$ are not defined a priori as a direct sum of pre-existing mixed Hodge structures on the cohomology of each sector.
If one looks for a polarization of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$, it is natural to use the orbifold Poincaré duality defined by Chen and Ruan in [3] since it shares many properties with the duality of $\mathcal{D}$-modules studied in [29] (this is highlighted in Appendix B.1 and Appendix B.2). In Theorem 5.5 we get such a polarization by considering a suitable normalization/Tate-twist of this orbifold Poincaré duality. Of course, it is possible to define different polarizations, depending on the specific Poincaré-Lefschetz conjugations involved: see Section 5.2 where another polarization is discussed, close to the one constructed by Fernandez in [15, Theorem 5.3] (in this last case we have to assume moreover that the Hodge-Tate condition (4) is satisfied).
Throughout this text, we will test our results on weighted projective spaces (the corresponding polytopes are reduced simplices in the sense of [4]), see for instance [6], [12], [23], [24] (mirror partners on the singularity side appear already in [14]). This is a nice setting to work with: quick computations can be carried out and, remarkably, almost everything can be deduced from the weight of the corresponding simplex. For instance, we get a very simple and effective formula for the Hodge numbers of the corresponding mixed Hodge structures $\operatorname{MHS}_{\operatorname{orb}}$ and $\operatorname{MHS}_{f}$, depending only on the ages of the sectors (see Corollary 3.17). We get also nice formulas for the primitive subspaces and their Hodge decomposition and this allows us to get a polarization on the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ rather inexpensively, see Proposition 5.6.
These notes are organized as follows: in Section 2, we describe and compare the three different settings considered in this paper; in Section 3, we define a mixed Hodge structure on the orbifold cohomology and we give a formula for its Hodge numbers; we analyze the Hodge-Tate condition in Section 4 and we discuss the existence of a polarization via the orbifold Poincaré duality in Section 5. The definitions about polarized mixed Hodge structures used in this text are recalled in Appendix A. The link between the polarization constructed in this paper and the duality of $\mathcal{D}$-modules is emphasized in Appendix B.1 and Appendix B.2.
2 The settings
This section is expository in nature.
2.1 The graded Milnor ring
Let $f(x)=\sum_{m}a_{m}x^{m}$, where $m=(m_{1},\ldots,m_{n})\in\mathbb{Z}^{n}$ and $x^{m}=x_{1}^{m_{1}}\ldots x_{n}^{m_{n}}$, be a Laurent polynomial defined on $(\mathbb{C}^{*})^{n}$. Let $\operatorname{supp}(f)=\{m,a_{m}\neq 0\}$. The Newton polytope $P$ of $f$ is the convex hull of $\operatorname{supp}(f)$ in $\mathbb{R}^{n}$. We will say that $f$ is convenient if $P$ contains the origin as an interior point and that it is nondegenerate if, for each face $\Delta$ of $P$,
$x_{1}\frac{\partial f_{\Delta}}{\partial x_{1}},\ldots,x_{n}\frac{\partial f_{\Delta}}{\partial x_{n}}$
do not vanish simultaneously on $(\mathbb{C}^{*})^{n}$, where $f_{\Delta}(x):=\sum_{m\in\Delta}a_{m}x^{m}$.
Let $\mathcal{A}_{f}:=\mathcal{B}/\mathcal{L}$ where $\mathcal{B}:=\mathbb{C}[u_{1},u_{1}^{-1},\ldots,u_{n},u_{n}^{-1}]$ and $\mathcal{L}:=(u_{1}\frac{\partial f}{\partial u_{1}},\ldots,u_{n}\frac{\partial f}{\partial u_{n}})$ is the ideal generated by the partial derivative $u_{1}\frac{\partial f}{\partial u_{1}},\ldots,u_{n}\frac{\partial f}{\partial u_{n}}$ of $f$. We define an increasing filtration $\mathcal{N}_{\bullet}$ on $\mathcal{B}$, indexed by $\mathbb{Q}$, by setting
$\mathcal{N}_{\alpha}\mathcal{B}:=\{g\in\mathcal{B},\ \operatorname{supp}(g)\in\nu^{-1}(]-\infty;\alpha])\}$
where $\nu:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the Newton function of the Newton polytope $P$ of $f$ (the function which takes the value $1$ at the vertices of $P$ and which is linear on each cone generated by the faces of $P$) and $\operatorname{supp}(g)=\{m\in\mathbb{Z}^{n},\ a_{m}\neq 0\}$ if $g=\sum_{m\in\mathbb{Z}^{n}}a_{m}u^{m}\in\mathcal{B}$. By projection, the Newton filtration $\mathcal{N}_{\bullet}$ on $\mathcal{B}$ induces a filtration, also denoted by $\mathcal{N}_{\bullet}$, on $\mathcal{A}_{f}$. We will denote by $A^{*}$ the graded ring
$A^{*}:=\operatorname{gr}^{\mathcal{N}}_{*}\mathcal{A}_{f}$
and we call it the graded Milnor ring. Note that the multiplication by $f$ on $\mathcal{A}_{f}$ induces a map $[f]$ of degree one on $A^{*}$. This is Kouchnirenko’s setting [22].
2.2 Orbifold cohomology and simplicial polytopes
We now briefly recall the setting of Borisov, Chen and Smith [2]. Let $P$ be a simplicial lattice polytope in $\mathbb{R}^{n}$, containing the origin as an interior point. We will denote by $\mathcal{V}(P)$ the set of its vertices. Let
  • $\mathcal{X}_{P}$ be the Deligne-Mumford stack associated by [2, Section 3] with the stacky fan $\mathbf{\Sigma}_{P}:=(\mathbb{Z}^{n},\Sigma_{P},\mathcal{V}(P))$, where $\Sigma_{P}$ is the fan whose cones are the cones over the faces of $P$,
  • $I_{\mathcal{X}_{P}}=\coprod_{\ell\in F}\mathcal{X_{\ell}}$ be the decomposition into connected components of the inertia orbifold of $\mathcal{X}_{P}$,
  • $H_{\operatorname{orb}}^{2\alpha}(\mathcal{X}_{P},\mathbb{Q}):=\oplus_{\ell\in F}H^{2(\alpha-i_{\ell})}(X_{\ell},\mathbb{Q})$ be the orbifold cohomology groups of $\mathcal{X}_{P}$ where $X_{\ell}$ denotes the underlying space of the sector $\mathcal{X}_{\ell}$ and $i_{\ell}:=\operatorname{age}(\mathcal{X}_{\ell})$ denotes its age.
We will write $\mathcal{X}$ instead of $\mathcal{X}_{P}$ and $H_{\operatorname{orb}}^{2\alpha}(\mathcal{X})$ instead of $H_{\operatorname{orb}}^{2\alpha}(\mathcal{X}_{P},\mathbb{C})$. We will consider the decomposition (of $\mathbb{C}$-vector spaces)
$H_{\operatorname{orb}}^{2*}(\mathcal{X})=H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}\oplus H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$ (5)
where
$H_{\operatorname{orb}}^{2\alpha}(\mathcal{X})_{0}:=\bigoplus_{\ell\in F,\ i_{\ell}\in\mathbb{Z}}H^{2(\alpha-i_{\ell})}(X_{\ell},\mathbb{C})\ \mbox{and}\ H_{\operatorname{orb}}^{2\alpha}(\mathcal{X})_{\neq 0}:=\bigoplus_{\ell\in F,\ i_{\ell}\notin\mathbb{Z}}H^{2(\alpha-i_{\ell})}(X_{\ell},\mathbb{C}).$
We will also make a repeated use of the equality
$n_{\ell}+i_{\ell}+i_{\ell^{-1}}=n$ (6)
given by [3, Lemma 3.2.1], where $n_{\ell}:=\dim X_{\ell}$ and $\ell^{-1}$ denotes the image of $\ell$ under the involution of $F$ induced by the involution $I$ of $I_{\mathcal{X}}$ defined by $I((p,(g))=(p,(g^{-1}))$ where $(g)$ denotes the conjugacy class of $g$ in $G_{p}$ (the local group at $p$), see [3, Remark 3.1.4].
In our toric setting, we have the following description:
  • for $\sigma$ a cone in the fan $\Sigma_{P}$, let $\operatorname{Box}(\sigma)$ be the set of the elements $v\in N$ such that
    $v=\sum_{\rho_{i}\subseteq\sigma}q_{i}b_{i}\ \mbox{for some}\ 0<q_{i}<1$
    where $\rho_{i}$ is the ray generated by the vertex $b_{i}$ of $P$. Let $\operatorname{Box}(\mathbf{\Sigma}_{P})$ be the union of $\operatorname{Box}(\sigma)$ for all cones $\sigma\in\Sigma_{P}$. By [2, Proposition 4.7], the sectors are parametrized by $v\in\operatorname{Box}(\mathbf{\Sigma}_{P})$.
  • For $v\in\operatorname{Box}(\mathbf{\Sigma}_{P})$, $X_{v}$ is the (simplicial) toric variety of the quotient fan $\Sigma_{P}/\sigma(v)$, where $\sigma(v)$ is the smallest cone of $\Sigma_{P}$ containing $v$ and $n_{v}:=\dim X_{v}=n-\dim\sigma(v)$.
  • If $v\in\operatorname{Box}(\mathbf{\Sigma}_{P})$, the age of the sector $X_{v}$ is $i_{v}=\nu(v)$ where $\nu$ is the Newton function of $P$ defined in Section 2.1, see [2, Remark 5.4].
  • For $v=\sum_{\rho_{i}\subseteq\sigma(v)}q_{i}b_{i}\in\operatorname{Box}(\sigma(v))$, let us define $v^{-1}:=\sum_{\rho_{i}\subseteq\sigma(v)}(1-q_{i})b_{i}$. Then $v^{-1}\in\operatorname{Box}(\sigma(v))$ and
    $\nu(v)+\nu(v^{-1})=\dim\sigma(v)=n-n_{v}$
    which is precisely equality (6) above. The involution $I$ on the labels is defined by $I(v)=v^{-1}$.
2.3 Sabbah’s mixed Hodge structure
In this section, we recall Sabbah’s setting [28], [29]. Let $f$ be a convenient and nondegenerate Laurent polynomial on $(\mathbb{C}^{*})^{n}$, where $n\geq 2$. We will denote by $G$ the localized Laplace transform of its Gauss-Manin system and by $G_{0}$ its Brieskorn module:
$G=\Omega^{n}(U)[\theta,\theta^{-1}]/(\theta d-df\wedge)\Omega^{n-1}(U)[\theta,\theta^{-1}]$
and $G_{0}$ is the image of $\Omega^{n}(U)[\theta]$ in $G$. Since by [22] $f$ has only isolated singularities, we have
$G_{0}=\Omega^{n}(U)[\theta]/(\theta d-df\wedge)\Omega^{n-1}(U)[\theta]$
and
$G_{0}/\theta G_{0}\cong\Omega^{n}(U)/df\wedge\Omega^{n-1}(U),$
the latter being a finite dimensional vector space of dimension $\mu$, the global Milnor number of $f$. Actually, $G_{0}$ is a lattice in $G$, i.e $G_{0}$ is free over $\mathbb{C}[\theta]$ and $\mathbb{C}[\theta,\theta^{-1}]\otimes_{\mathbb{C}[\theta]}G_{0}=G$: this is shown in [13, Section 4], using the main properties of convenient and nondegenerate functions established in [22]. More generally, this is a characteristic property of tame functions, for which Sabbah’s mixed Hodge structures can be constructed [29, Section 13].
Let $\tau:=\theta^{-1}$. Then $G$ is a free $\mathbb{C}[\tau,\tau^{-1}]$-module, equipped with a derivation $\partial_{\tau}$ (induced from the formula $\partial_{\tau}[\omega]=[-f\omega]$ if $\omega\in\Omega^{n}(U)$ where the brackets denote the class in $G$). Let $V_{\bullet}G$ be the Kashiwara-Malgrange filtration of $G$ along $\tau=0$ as in [29, Section 1]. This is an increasing filtration, indexed by $\mathbb{Q}$, which satisfies the following properties:
  • for every $\alpha\in\mathbb{Q}$, $V_{\alpha}G$ is a free $\mathbb{C}[\tau]$-module of rank $\mu$,
  • $\tau V_{\alpha}G\subset V_{\alpha-1}G$ and $\partial_{\tau}V_{\alpha}G\subset V_{\alpha+1}G$,
  • $\tau\partial_{\tau}+\alpha$ is nilpotent on $\operatorname{gr}_{\alpha}^{V}G$.
We define, for $\alpha\in\mathbb{Q}$,
$H_{\alpha}:=\operatorname{gr}_{\alpha}^{V}G\ \mbox{and}\ H:=\oplus_{\alpha\in[0,1[}H_{\alpha}$
and we write $H=H_{\neq 0}\oplus H_{0}$, where $H_{\neq 0}:=\oplus_{\alpha\in]0,1[}H_{\alpha}$. The $\mathbb{C}$-vector space $H$ (of dimension $\mu$) is equipped with a nilpotent endomorphism $N$ induced by $(\tau\partial_{\tau}+\alpha)$ on $H_{\alpha}$.
The Hodge filtration. Let $G_{\bullet}$ be the increasing filtration of $G$ defined by $G_{p}:=\theta^{-p}G_{0}=\tau^{p}G_{0}$ and
$G_{p}H_{\alpha}:=G_{p}\cap V_{\alpha}G/G_{p}\cap V_{<\alpha}G$
for $\alpha\in[0,1[$ and $p\in\mathbb{Z}$. The Hodge filtration on $H$ is the decreasing filtration $F^{\bullet}$ defined by
$F^{p}H_{\alpha}:=G_{n-1-p}H_{\alpha}\ \mbox{if}\ \alpha\in]0,1[\ \mbox{and}\ F^{p}H_{0}:=G_{n-p}H_{0}$ (7)
for $p\in\mathbb{Z}$.
The real structure and the weight filtration. We have on $H$ a real structure coming from the identification
$H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{n}(U,f^{-1}(t);\mathbb{C})=H^{n}(U,f^{-1}(t);\mathbb{R})\otimes_{\mathbb{R}}\mathbb{C}$
(for $|t|>>0$): in this setting, $H_{\alpha}$ corresponds to the generalized eigenspace of the monodromy at infinity associated with the eigenvalue $\exp(2i\pi\alpha)$ and the unipotent part of this monodromy is equal to $\exp(2i\pi N)$. We will write $H=H^{\mathbb{R}}\otimes\mathbb{C}$. We have the decomposition $H^{\mathbb{R}}=H_{0}^{\mathbb{R}}\oplus H_{\neq 0}^{\mathbb{R}}$ and the weight filtration $W_{\bullet}$ on $H^{\mathbb{R}}$ is the increasing weight filtration of $2i\pi N$ centered at n1n-1 on $H_{\neq 0}^{\mathbb{R}}$ and centered at $n$ on $H_{0}^{\mathbb{R}}$.
Sabbah’s mixed Hodge structures. By [29, Theorem 13.1], the triple $\operatorname{MHS}_{f}:=(H,F^{\bullet},W_{\bullet})$ is a mixed Hodge structure.
Remark 2.1
Deligne’s canonical subspaces $I^{p,q}$ (see [7]) provide the splittings $F^{p}H=\bigoplus_{i\geq p}I^{i,j}$ and $W_{m}=\bigoplus_{i+j\leq m}I^{i,j}$. Using the decomposition $F^{p}H=(\oplus_{j}I^{p,j})\bigoplus F^{p+1}H$, we get an isomorphism $H\cong\operatorname{gr}_{F}H$ such that $F^{p}H\cong\bigoplus_{i\geq p}\operatorname{gr}_{F}^{i}H$ and, since $N(I^{p,q})\subset I^{p-1,q-1}$, a commutative diagram
[…]
where $[N]$ is induced by $N]:\operatorname{gr}_{F}^{i}H\rightarrow\operatorname{gr}_{F}^{i-1}H$: in particular, $N$ and $[N]$ have the same Jordan normal form (see [34, Theorem 7.1]).
2.4 Isomorphisms
We recall the mirror theorem of Borisov, Chen and Smith [2].
Proposition 2.2 ([2])
Let $P$ be a simplicial lattice polytope in $\mathbb{R}^{n}$ containing the origin as an interior point and let $f$ be the Laurent polynomial defined on $(\mathbb{C}^{*})^{n}$ by
$f(u)=\sum_{b\in\mathcal{V}(P)}u^{b}$ (8)
where $\mathcal{V}(P)$ denotes the set of the vertices of its Newton polytope $P$. Then, we have an isomorphism
$\varphi:H_{\operatorname{orb}}^{2*}(\mathcal{X})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}A^{*}$ (9)
of $\mathbb{Q}$-graded rings where $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ is the orbifold cohomology of $P$ (equipped with the orbifold cup-product) defined in Section 2.2 and $A^{*}$ is the graded Milnor ring of $f$ defined in Section 2.1.
Proof.  Note first that the assumptions of Proposition 2.2 imply that $f$ is convenient and nondegenerate. It remains to identify the graded ring appearing in [2, Theorem 1.1] with $A^{*}$. Recall that if $\partial:\mathcal{C}\rightarrow\mathcal{A}$ is a strict morphism of filtered module (i.e $\partial(\mathcal{C})\cap\mathcal{A}_{q}=\partial(\mathcal{C}_{q})$) we have $\operatorname{gr}(\mathcal{A}/\partial\mathcal{C})\cong\operatorname{gr}(\mathcal{A})/\operatorname{gr}(\partial)\operatorname{gr}(\mathcal{C})$, see for instance [22, Lemme 4.4]. We get the expected identification by applying this observation to the morphism
$\partial:\mathcal{A}^{n}\longrightarrow\mathcal{A}$
defined by
$\partial(g_{1},\ldots,g_{n})=g_{1}x_{1}\frac{\partial f}{\partial x_{1}}+\ldots+g_{n}x_{n}\frac{\partial f}{\partial x_{n}},$
which is strict for the following choice of the filtration in $\mathcal{A}^{n}$: the Newton weight of $(0,\ldots,g_{i},\ldots,0)$ in is equal to $\nu(g_{i})-1$, see [22, Theorem 4.1].   $\Box$
Remark 2.3
Proposition 2.2 provides an isomorphism of rings, and this depends on the special form of $f$, from which we also get $\omega:=\varphi^{-1}([f])\in H^{2}(X_{0})\subset H_{\operatorname{orb}}^{2}(\mathcal{X})$ where $X_{0}$ denotes the untwisted sector (recall that $f]:A^{*}\longrightarrow A^{*}$ denotes the nilpotent endomorphism induced by the multiplication by $f$ on $\mathcal{A}_{f}$).
Last, we connect Sabbah’s mixed Hodge structures and graded Milnor rings. The following result is based on the description of Sabbah’s Hodge filtration with the help of the Brieskorn lattice and a Kashiwara-Malgrange $V$-filtration and the identification of the latter with a suitably normalized Newton filtration provided by [13, Theorem 4.5].
Proposition 2.4 ([11], [13], [29], [30])
Let $f$ be a convenient and nondegenerate Laurent polynomial and let $F^{\bullet}$ be the Hodge filtration of its Sabbah’s mixed Hodge structure. Then, for $p\in\mathbb{Z}$ we have an isomorphism,
$\operatorname{gr}_{F}^{p}H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\bigoplus_{n-1-p<\alpha\leq n-p}A^{\alpha}$
In particular, the graded vector space $\operatorname{gr}_{F}H=\oplus_{p}\operatorname{gr}_{F}^{p}H$ is isomorphic to its graded Milnor ring $A^{*}$.   $\Box$
Last, the following well-known result is useful in order to describe $N$:
Proposition 2.5 ([29], [34], [41])
The maps $[f]$ and NN have the same Jordan normal form.
Proof.  We follow [34, Theorem 7.1]. For $\beta\in]0,1[$, we have the commutative diagram
$A^{\beta+n-1-p}$$A^{\beta+n-p}$$\operatorname{gr}_{F}^{p}H_{\beta}$$\operatorname{gr}_{F}^{p-1}H_{\beta}$$\scriptstyle{[f]}$$\scriptstyle{\tau^{(n-1-p)}}$$\scriptstyle{\tau^{n-p}}$$\scriptstyle{[N]}$
thanks to Proposition 2.4 (we use here the fact that $N]=-\tau\partial_{\tau}\mod F^{p}$ and $\partial_{\tau}[\omega]=[-f\omega]$), hence $[f]$ and $[N]$ have the same Jordan normal form. By Remark 2.1, NN and $[N]$ have the same Jordan normal form and the assertion follows. Analogous argument for $H_{0}$.   $\Box$
2.5 Résumé
Let $P$ be a simplicial polytope containing the origin as an interior point and let $f$ be the Laurent polynomial defined by $f(u)=\sum_{b\in\mathcal{V}(P)}u^{b}$. Then $f$ is a nondegenerate and convenient and we have the isomorphisms
$F^{p}H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\bigoplus_{i\geq p}\operatorname{gr}_{F}^{i}H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\bigoplus_{\alpha\leq n-p}A^{\alpha}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\bigoplus_{\alpha\leq n-p}H^{2\alpha}_{\operatorname{orb}}(\mathcal{X})$ (10)
and the isomorphisms of complex vector spaces
$H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{gr}_{F}H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}A^{*}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{2*}_{\operatorname{orb}}(\mathcal{X}).$ (11)
Under these isomorphisms, $N$, $[N]$ and $[f]$ correspond respectively yo $[N]$, $[f]$ and $L_{\omega}$ where $L_{\omega}$ denotes the nilpotent endomorphism of $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ defined by
$L_{\omega}(a):=\omega\cup_{\operatorname{orb}}a$ (12)
and $\omega:=\varphi^{-1}([f])\in H^{2}(X_{0})$. We will denote by
$\psi:H\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{2*}_{\operatorname{orb}}(\mathcal{X})$ (13)
the composition of these isomorphisms.
Remark 2.6
[…]
Remark 2.7
The previous isomorphisms are helpful in order to understand better the spectrum at infinity and the monodromy at infinity of a Laurent polynomial $f$.
  1. 1.
    The spectrum at infinity of $f$ is defined by $\operatorname{Spec}_{f}(z)=\sum_{\alpha\in\mathbb{Q}}\dim A^{\alpha}z^{\alpha}$ and one can use the isomorphism (9) and the description of the orbifold cohomology given in Section 2.2 to compute this sequence.
  2. 2.
    By Proposition 2.5, NN and $[f]$ have the same Jordan normal form. The isomorphisms (11) gives information about this Jordan structure.
Example 2.8
(Weighted projective spaces) We follow the terminology of [4]: we will say that $P$ is a simplex in $\mathbb{R}^{n}$ if $P=\operatorname{conv}(v_{0},\cdots,v_{n})$ where $v_{i}\in\mathbb{Z}^{n}$ for $i=0,\ldots n$ and if it contains the origin as an interior point. The weight of the simplex $P$ is the tuple $(q_{0},\cdots,q_{n})$ where
$q_{i}:=|\det(v_{0},\cdots,\widehat{v_{i}},\cdots,v_{n})|$
for $i=0,\cdots,n$ (we will always assume that the $q_{i}$’s are arranged in increasing order). The simplex $P$ is reduced if $\gcd(q_{0},\cdots,q_{n})=1$ (this condition is already used in [14, p. 2]). The corresponding orbifold $\mathcal{X}_{P}$ is a weighted projective space.
  • Using the isomorphisms (11) and the description of the corresponding Sabbah’s mixed Hodge structure given in [14], it is possible to give informations about $H^{2*}_{\operatorname{orb}}(\mathcal{X}_{P})$ (see also [6], [12], [24]…). Let
    $F:=\left\{\frac{j}{q_{i}}|\,0\leq j\leq q_{i}-1,\ 0\leq i\leq n\right\}$
    and let $f_{1},\cdots,f_{k}$ be the elements of $F$ arranged by increasing order. Define
    $S_{f_{i}}:=\{j|\ q_{j}f_{i}\in\mathbb{Z}\}\subset\{0,\cdots,n\}\ \mbox{and}\ d_{i}:=\operatorname{Card}S_{f_{i}}.$
    Then, the sectors are labelled by $F$ and the age of the sector $X_{f_{i}}$ for $f_{i}\in F$ is
    $i_{f_{i}}=d_{1}+\cdots+d_{i-1}-\mu f_{i}\ \mbox{and}\ n_{f_{i}}:=\dim X_{f_{i}}=d_{i}-1$ (14)
    for $i=2,\cdots,k$, where $\mu:=q_{0}+\ldots+q_{n}$.
  • Conversely, each sector $X_{\ell}$ produces exactly one Jordan block of size $n_{\ell}+1$ of the nilpotent endomorphism $N$ of $H$ since, by [21], $\dim H^{2i}(X_{\ell})=1$ and $H^{2i+1}(X_{\ell})=0$ for all $\ell\in F$. Moreover, the spectrum at infinity is equal the disjoint union over $\ell\in F$ of the sets
    $\{i_{\ell},i_{\ell}+1,\ldots,i_{\ell}+n_{\ell}\}.$
    This gives another interpretation of the results obtained in [14].
Note that the ages of the sectors are all integral numbers if and only $q_{i}$ divides $\mu$ for $i=0,\ldots,n$, that is if and only if $P$ is a reflexive polytope.
3 A mixed Hodge structure on orbifold cohomology
Let $P$ be a simplicial lattice polytope containing the origin as an interior point and let $f$ be the convenient and nondegenerate Laurent polynomial defined on $(\mathbb{C}^{*})^{n}$ by
$f(u)=\sum_{b\in\mathcal{V}(P)}u^{b}$
as in (8). In this section, we construct mixed Hodge structures on $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ corresponding to the mixed Hodge structures $\operatorname{MHS}_{f}$.
3.1 Filtrations
The definition of the following Hodge and weight filtrations was originally suggested by the description given in [14] of the Sabbah’s mixed Hodge structures associated with Laurent polynomials whose Newton polytopes are reduced simplices. In a more combinatorial setting, analogous filtrations are defined by Wang in [42].
3.1.1 The Hodge filtration
We define, on each sector,
$F^{p}H^{2*}(X_{\ell},\mathbb{C}):=\oplus_{i\geq p}H^{2(n-i-i_{\ell})}(X_{\ell},\mathbb{C})\ \mbox{if}\ i_{\ell}\in\mathbb{Z},$ (15)
and
$F^{p}H^{2*}(X_{\ell},\mathbb{C}):=\oplus_{i\geq p}H^{2(n-1-i-[i_{\ell}])}(X_{\ell},\mathbb{C})\ \mbox{if}\ i_{\ell}\notin\mathbb{Z}$ (16)
for $p\in\mathbb{Z}$.
Definition 3.1
The Hodge filtration on $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is the decreasing filtration defined by
$F^{p}H_{\operatorname{orb}}^{2*}(\mathcal{X}):=\oplus_{\ell\in F}F^{p}H^{2*}(X_{\ell},\mathbb{C})$ (17)
for $p\in\mathbb{Z}$.
This Hodge filtration has a very simple description:
Proposition 3.2
Let $p\in\mathbb{Z}$. Then,
$F^{n-p}H_{\operatorname{orb}}^{2*}(\mathcal{X})=\bigoplus_{\alpha\leq p}H_{\operatorname{orb}}^{2\alpha}(\mathcal{X}).$ (18)
In particular, $F^{0}H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{C})=H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{C})$ and $F^{n+1}H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{C})=0$.
Proof.  We have
$\bigoplus_{p-1<\alpha\leq p}H_{\operatorname{orb}}^{2\alpha}(\mathcal{X})=\oplus_{i_{\ell}\in\mathbb{Z}}H^{2(p-i_{\ell})}(X_{\ell})\bigoplus\oplus_{i_{\ell}\notin\mathbb{Z}}H^{2(p-1-[i_{\ell}])}(X_{\ell})$
and (18) follows.   $\Box$
3.1.2 The weight filtration
We define, on each sector,
$W_{m}H^{2*}(X_{\ell},\mathbb{C}):=\bigoplus_{n+n_{\ell}-m\leq 2j}H^{2j}(X_{\ell},\mathbb{C})\ \mbox{if}\ i_{\ell}\in\mathbb{Z}$ (19)
and
$W_{m}H^{2*}(X_{\ell},\mathbb{C}):=\bigoplus_{n-1+n_{\ell}-m\leq 2j}H^{2j}(X_{\ell},\mathbb{C})\ \mbox{if}\ i_{\ell}\notin\mathbb{Z}$ (20)
for $m\in\mathbb{Z}$.
Definition 3.3
The weight filtration on $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is the increasing filtration defined by
$W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})=\oplus_{\ell\in F}W_{m}H^{2*}(X_{\ell})$
for $m\in\mathbb{Z}$.
Recall the nilpotent endomorphism $L_{\omega}$ of $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ defined by (12).
Proposition 3.4
The weight filtration $W_{\bullet}$ is the weight filtration of $L_{\omega}$, centered at $n$ on $H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}$ and centered at n1n-1 on $H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$.
Proof.  Note first that $W_{2n}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}=H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}\ \mbox{and}\ W_{-1}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}=0$ because $n_{\ell}\leq n$ if $i_{\ell}\in\mathbb{Z}$ and $W_{2(n-1)}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}\ \mbox{and}\ W_{-1}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=0$ because $n_{\ell}\leq n-1$ if $i_{\ell}\notin\mathbb{Z}$. Using the characteristic properties of the weight filtration of a nilpotent endomorphism, it is enough to check that
  • $L_{\omega}^{n+1}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}=0$, $L_{\omega}(W_{k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0})\subset W_{k-2}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}$ and
    $L_{\omega}^{k}:\operatorname{gr}^{W}_{n+k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{gr}^{W}_{n-k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}$ (21)
    for $k\leq n/2$,
  • $L_{\omega}^{n}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=0$, $L_{\omega}(W_{k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0})\subset W_{k-2}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$ and
    $L_{\omega}^{k}:\operatorname{gr}^{W}_{n-1+k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{gr}^{W}_{n-1-k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$ (22)
    for $k\leq(n-1)/2$.
In both cases, the two first properties follow immediately from the definition of $L_{\omega}$. Let us show the third on for $H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}$. Recall that $L_{\omega}$ preserves each sector (see Remark 2.6). By the hard Lefschetz property for the toric simplicial variety $X_{\ell}$ (see for instance [5, Theorem 12.5.8] or [15, Theorem 2.2] and the references therein; this is easily obtained when $P$ is a reduced simplex, see Remark 5.7 below), we have
$L_{\omega}^{k}:H^{n_{\ell}-k}(X_{\ell})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}H^{n_{\ell}+k}(X_{\ell})$
and therefore
$L_{\omega}^{k}:\operatorname{gr}^{W}_{n+k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{gr}^{W}_{n-k}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}$
since $\operatorname{gr}^{W}_{n+k}H^{2*}_{\operatorname{orb}}(\mathcal{X})_{0}=\oplus_{i_{\ell}\in\mathbb{Z}}H^{n_{\ell}-k}(X_{\ell})$. The assertion about $H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$ is shown in the same way.   $\Box$
3.1.3 Comparison with Sabbah’s filtrations
We compare filtrations defined above on the orbifold cohomology with the filtrations of the mixed Hodge structure $\operatorname{MHS}_{f}$.
Proposition 3.5
The Hodge filtration (resp. the weight filtration) on the orbifold cohomology defined in Definition 3.1 (resp. Definition 3.3) is the image, under the isomorphism (13), of the Hodge filtration (resp. of the weight filtration), defined on the complex vector space $H$, of the mixed Hodge structure $\operatorname{MHS}_{f}$.
Proof.  By (10) and Proposition 3.2, the Hodge filtration on the orbifold cohomology is the image of Sabbah’s Hodge filtration under the isomorphism $\psi$. The assertion about the weight filtration follows from (13) and Proposition 3.4.   $\Box$
3.2 Conjugations
The last ingredient that we need in order to define a mixed Hodge structure is a conjugation on $H_{\operatorname{orb}}^{2*}(\mathcal{X})$. It turns out that the correct ones to consider are the following ones (see the proof of Theorem 3.10 below):
Definition 3.6
We will call Poincaré-Lefschetz conjugation on $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ any conjugation satisfying
$\overline{H^{2i}(X_{\ell})}\subset\oplus_{j\geq i}H^{2j}(X_{\ell^{-1}})$ (23)
for each sector $X_{\ell}$.
A Poincaré-Lefschetz conjugation defines a real structure on $W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})$:
Lemma 3.7
Let $W_{\bullet}$ be the weight filtration on the orbifold cohomology defined in Definition 3.3. Then, for any Poincaré-Lefschetz conjugation,
$\overline{W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})}=W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X}).$
In particular,
$W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})=(W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})\cap H_{\operatorname{orb},\mathbb{R}}^{2*}(\mathcal{X}))\otimes\mathbb{C}$
where $H_{\operatorname{orb},\mathbb{R}}^{2*}(\mathcal{X})$ denotes the real form defined by the Poincaré-Lefschetz conjugation involved.
Proof.  By the very definition of Poincaré-Lefschetz conjugations, and since $n_{\ell}=n_{\ell^{-1}}$, we have
$\overline{W_{m}H^{2*}(X_{\ell})}\subset W_{m}H^{2*}(X_{\ell^{-1}})$
for all $\ell\in F$. The assertion follows.   $\Box$
Remark 3.8
We will mainly consider the following Poincaré-Lefschetz conjugations:
  • the Poincaré-Lefschetz conjugation defined by
    $\overline{a}:=(-1)^{i}I^{*}(c(a))\in H^{2i}(X_{\ell^{-1}},\mathbb{C})$ (24)
    if $a\in H^{2i}(X_{\ell},\mathbb{C})$.
  • The Poincaré-Lefschetz conjugation defined by
    $\overline{a}:=I^{*}(c(a))\in H^{2i}(X_{\ell^{-1}},\mathbb{C})$ (25)
    if $a\in H^{2i}(X_{\ell},\mathbb{C})$.
Here, $c$ denotes the standard conjugation on $H^{*}(X_{\ell^{-1}},\mathbb{C})=H^{*}(X_{\ell^{-1}},\mathbb{R})\otimes\mathbb{C}$ and $I$ is the complexification of the involution $I:H^{*}(X_{\ell^{-1}},\mathbb{R})\rightarrow H^{*}(X_{\ell},\mathbb{R})$ induced by the involution $I$ considered in Section 2.2.
The following lemma motivates the definition of the conjugations (24) and (25).
Lemma 3.9
Assume that $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ is equipped with the conjugation (24) (resp. (25)). Then,
  1. 1.
    $I^{*}(L_{\omega}(a))=L_{\omega}(I^{*}(a))$,
  2. 2.
    $\overline{L_{\omega}(a)}=-L_{\omega}(\overline{a})$ (resp. $\overline{L_{\omega}(a)}=L_{\omega}(\overline{a})$),
  3. 3.
    $\overline{2i\pi L_{\omega}(a)}=2i\pi L_{\omega}(\overline{a})$ (resp. $\overline{L_{\omega}(a)}=L_{\omega}(\overline{a})$).
Proof.  The first equality follows from Remark 2.6. We show the remaining statements when $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ is equipped with the conjugation (24). Since $L_{\omega}$ is real with respect to the standard conjugation $c$, we get
$I^{*}(\overline{L_{\omega}(a)})=(-1)^{i+1}c(L_{\omega}(a))=(-1)^{i+1}L_{\omega}(c(a))=-L_{\omega}(I^{*}(\overline{a}))=-I^{*}(L_{\omega}(\overline{a})).$
The second and the third assertions follow.   $\Box$
Let us note that conjugation (25) appears implicitly in [42, Proposition 4.7], but it does not match with the conjugations appearing in Sabbah’s mixed Hodge structures, for which $2i\pi L_{\omega}(=2i\pi N)$ is real.
3.3 A mixed Hodge structure
We now define a mixed Hodge structure on $H_{\operatorname{orb}}^{2*}(\mathcal{X})$.
Theorem 3.10
Assume that $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ is equipped with a Poincaré-Lefschetz conjugation. Then, the tuple  [^1] Strictly speaking, $\operatorname{MHS}_{\operatorname{orb}}$ should be indexed by the Poincaré-Lefschetz conjugation involved, but there will be no risk of confusion in what follows.
$\operatorname{MHS}_{\operatorname{orb}}:=(H_{\operatorname{orb}}^{2*}(\mathcal{X}),H_{\operatorname{orb},\mathbb{R}}^{2*}(\mathcal{X}),F^{\bullet},W_{\bullet})$
is a mixed Hodge structure.
Proof.  We check that
$\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})=F^{p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})\oplus\overline{F^{m+1-p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})}$ (26)
for $m,p\in\mathbb{Z}$. It is enough to show (26) for each summand of the direct sum decomposition (5) and we first begin with $H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$. Recall that
$\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}H^{n-1+n_{\ell}-m}(X_{\ell}).$ (27)
We have
$W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}(H^{n-1+n_{\ell}-m}(X_{\ell})\oplus H^{n+n_{\ell}-m}(X_{\ell})\oplus\ldots\oplus H^{2n_{\ell}}(X_{\ell}))$
and
$F^{p}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}(H^{2(n-1-p-[i_{\ell}])}(X_{\ell})\oplus H^{2(n-2-p-[i_{\ell}])}(X_{\ell})\oplus\ldots\oplus H^{0}(X_{\ell}))$
therefore
$F^{p}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}\cap W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}(H^{2(n-1-p-[i_{\ell}])}(X_{\ell})\oplus\ldots\oplus H^{n-1+n_{\ell}-m}(X_{\ell}))$
where the sum runs through the sectors such that $n-1+n_{\ell}-m\leq 2n-2-2p-2[i_{\ell}]$, in other words, by (6)), $m\geq 2p+[i_{\ell}]-[i_{\ell^{-1}}]$. Finally,
$F^{p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}H^{n-1+n_{\ell}-m}(X_{\ell})$
where the sum runs through the sectors such that $m\geq 2p+[i_{\ell}]-[i_{\ell^{-1}}]$ and
$F^{m+1-p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}H^{n-1+n_{\ell}-m}(X_{\ell})$
where the sum runs through the sectors such that $m\leq 2p-2-[i_{\ell}]+[i_{\ell^{-1}}]$. By the very definition Poincaré-Lefschetz conjugations, we deduce that
$\overline{F^{m+1-p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}}=\bigoplus_{i_{\ell}\notin\mathbb{Z}}H^{n-1+n_{\ell}-m}(X_{\ell})$
where the sum runs over the sectors such that $m\leq 2p-2+[i_{\ell}]-[i_{\ell^{-1}}]$. Last, there are no sectors satisfying $m=2p-1+[i_{\ell}]-[i_{\ell^{-1}}]$ in the definition of $\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$ since in this case $n-1+n_{\ell}-m=2n-2p-1+2[i_{\ell}]$ is odd and therefore $H^{n-1+n_{\ell}-m}(X_{\ell})=0$ in our toric setting. It follows from (27) that (26) holds for $H_{\operatorname{orb}}^{*}(\mathcal{X})_{\neq 0}$. Analogous calculations for $H_{\operatorname{orb}}^{2*}(\mathcal{X})_{0}$.   $\Box$
Remark 3.11
By the very definition of the filtrations, the direct sum decomposition (5) is a splitting of mixed Hodge structures, of respective weights $n$ and n1n-1.
3.4 Hodge numbers
We now compute the Hodge numbers $h^{p,q}:=\dim\operatorname{gr}_{F}^{p}\operatorname{gr}^{W}_{p+q}H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{C})$ of the mixed Hodge structures $\operatorname{MHS}_{\operatorname{orb}}$. According to the decomposition (5), we will write $h^{p,q}=h_{0}^{p,q}+h_{\neq 0}^{p,q}$.
Proposition 3.12
We have
$h_{\neq 0}^{p,q}=\sum_{i_{\ell}\notin\mathbb{Z},\ [i_{\ell}]-[i_{\ell^{-1}}]=q-p}\dim H^{2(n_{\ell}+[i_{\ell^{-1}}]-p)}(X_{\ell})$
and
$h_{0}^{p,q}=\sum_{i_{\ell}\in\mathbb{Z},\ i_{\ell}-i_{\ell^{-1}}=q-p}\dim H^{2(n_{\ell}+i_{\ell^{-1}}-p)}(X_{\ell})$
for $p,q\in\mathbb{Z}$.
Proof.  By the proof of Theorem 3.10, the Hodge decomposition of $\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}$ is
$\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}=\bigoplus_{p}F^{p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}\cap\overline{F^{m-p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}}$
where
$F^{p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}\cap\overline{F^{m-p}\operatorname{gr}^{W}_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})_{\neq 0}}=\bigoplus_{\ell}H^{2(n_{\ell}+[i_{\ell^{-1}}]-p)}(X_{\ell}),$
the sum running through the sectors $X_{\ell}$ such that $2p+[i_{\ell}]-[i_{\ell^{-1}}]=m$. This gives the first equality because the dimension of the subspace on the left of this last equality is precisely equal to $h^{p,m-p}_{\neq 0}$. The second equality is shown in the same way.   $\Box$
Remark 3.13
It follows from Proposition 3.12 that $h^{0,q}_{0}=0$ if $q>0$.
Recall that the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ is of Hodge-Tate type if $h^{p,q}=0$ for $p\neq q$ (see also Section 4 below).
Corollary 3.14
[…]
Proof.  The if part follows immediately from Proposition 3.12. Conversely, and again by Proposition 3.12, we have $h_{\neq 0}^{[i_{\ell^{-1}}],[i_{\ell}]}\geq\dim H^{2n_{\ell}}(X_{\ell})>0$ and $h_{0}^{i_{\ell^{-1}},i_{\ell}}\geq\dim H^{2n_{\ell}}(X_{\ell})>0$ therefore if there exists a label $\ell$ such that $i_{\ell}]\neq[i_{\ell^{-1}}]$ the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ is not of Hodge-Tate type.   $\Box$
Remark 3.15
By Proposition 3.5, the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{f}$ and $\operatorname{MHS}_{\operatorname{orb}}$ are the same. Therefore, the Hodge numbers of the mixed Hodge structures $\operatorname{MHS}_{f}$ satisfy the equalities of Proposition 3.12 and the conclusion of its corollaries. A formula for the Hodge numbers of the mixed Hodge structure $\operatorname{MHS}_{f}$ is also given in [42, Corollary 4.10] in terms of the $h$-vector of the Newton polytope $P$ of $f$.
Remark 3.16
Using Proposition 3.12, we easily get expected properties for the Hodge numbers of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ (see for instance [11, Section 2] for a summary of the properties of the Hodge numbers of $\operatorname{MHS}_{f}$). In particular,
  1. 1.
    $h^{p,q}_{0}=h^{q,p}_{0}$ and $h^{p,q}_{\neq 0}=h^{q,p}_{\neq 0}$,
  2. 2.
    $h^{n-p,n-q}_{0}=h^{p,q}_{0}$ and $h^{n-1-p,n-1-q}_{\neq 0}=h^{p,q}_{\neq 0}$.
The first point follows from the equality $\dim H^{2k}(X_{\ell})=\dim H^{2k}(X_{\ell^{-1}})$ and the second one from the Poincaré duality for the simplicial toric varieties $X_{\ell}$.
When $P$ is a reduced simplex, we get a formula depending only on the ages of the sectors and this is interesting because these ages are easily calculated from the weight of the simplex (see Example 2.8).
Corollary 3.17
Let $P$ be a reduced simplex. Then,
$h_{\neq 0}^{p,q}=\operatorname{Card}\{\ell\in F,\ [i_{\ell}]-[i_{\ell^{-1}}]=q-p,\ [i_{\ell}]\leq n-1-p,\ [i_{\ell^{-1}}]\leq p\}$
and
$h_{0}^{p,q}=\operatorname{Card}\{\ell\in F,\ i_{\ell}-i_{\ell^{-1}}=q-p,\ i_{\ell}\leq n-p,\ i_{\ell^{-1}}\leq p\}$
for $p,q\in\mathbb{Z}$.
Proof.  By [21], we have $\dim H^{2i}(X_{\ell})=1$ and $\dim H^{2i+1}(X_{\ell})=0$ for $i=0,\ldots,n_{\ell}$ and the formulas follow from Proposition 3.12 (the restrictions about the ages of the sectors follow from the inequalities $n_{\ell}\geq n_{\ell}+[i_{\ell^{-1}}]-p\geq 0$ and (6)).   $\Box$
Last, we give three examples: the first (resp. the third) one deals with the Hodge-Tate (resp. the reflexive) case and the second one illustrate the general case. We use Corollary 3.17 and Example 2.8.
Example 3.18
Let $f$ be the Laurent polynomial defined on $(\mathbb{C}^{*})^{5}$ by
$f(u_{1},u_{2},u_{3},u_{4},u_{5})=u_{1}+u_{2}+u_{3}+u_{4}+u_{5}+\frac{1}{u_{1}^{2}u_{2}^{2}u_{3}^{3}u_{4}^{3}u_{5}^{3}}.$
Its Milnor number is $\mu=14$. Its Newton polytope $P$ is the reduced simplex of weight $(1,2,2,3,3,3)$ and $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is the orbifold cohomology of the weighted projective space $\mathbb{P}(1,2,2,3,3,3)$ (see also [19]). The sectors are labelled by the set
$F=\{0,1/3,1/2,2/3\}$
with respective ages $i_{0}=0$, $i_{1/3}=4/3$, $i_{2/3}=5/3$, $i_{1/2}=2$ and dimensions $n_{0}=5$, $n_{1/3}=n_{2/3}=2$ and $n_{1/2}=1$. Since $(1/3)^{-1}=2/3$ and $(1/2)^{-1}=1/2$, the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ is of Hodge-Tate type and we get the Hodge diamond
[…]
$\begin{array}[]{ccccccccccc}&&&&&1&&&&&\\ &&&&0&&0&&&&\\ &&&0&&1&&0&&&\\ &&0&&0&&0&&0&&\\ &0&&0&&2&&0&&0&\\ 0&&0&&0&&0&&0&&0\\ &0&&0&&2&&0&&0&\\ &&0&&0&&0&&0&&\\ &&&0&&1&&0&&&\\ &&&&0&&0&&&&\\ &&&&&1&&&&&\\ \end{array}+\begin{array}[]{ccccccccccc}&&&&&0&&&&&\\ &&&&0&&0&&&&\\ &&&0&&2&&0&&&\\ &&0&&0&&0&&0&&\\ &0&&0&&2&&0&&0&\\ 0&&0&&0&&0&&0&&0\\ &0&&0&&2&&0&&0&\\ &&0&&0&&0&&0&&\\ &&&0&&0&&0&&&\\ &&&&0&&0&&&&\\ &&&&&0&&&&&\\ \end{array}$
for the mixed Hodge structures $\operatorname{MHS}_{\operatorname{orb}}$ and $\operatorname{MHS}_{f}$, where the sum refers to the direct sum decomposition (5).
Example 3.19
Let $f$ be the Laurent polynomial defined on $(\mathbb{C}^{*})^{5}$ by
$f(u_{1},u_{2},u_{3},u_{4},u_{5})=u_{1}+u_{2}+u_{3}+u_{4}+u_{5}+\frac{1}{u_{1}^{2}u_{2}^{3}u_{3}^{4}u_{4}^{7}u_{5}^{11}}.$
We have $\mu=28$. Its Newton polytope $P$ is the reduced simplex of weight $(1,2,3,4,7,11)$ and $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is the orbifold cohomology of the weighted projective space $\mathbb{P}(1,2,3,4,7,11)$. The sectors are labelled by the set
$F=\{0,1/11,1/7,2/11,1/4,3/11,2/7,1/3,4/11,3/7,5/11,1/2,6/11,4/7,7/11,2/3,5/7,$
8/11,3/4,9/11,6/7,10/11}8/11,3/4,9/11,6/7,10/11\}
with respective ages
0,38/11,3,32/11,2,26/11,3,8/3,31/11,2,25/11,2,30/11,3,24/11,7/3,2,0,38/11,3,32/11,2,26/11,3,8/3,31/11,2,25/11,2,30/11,3,24/11,7/3,2,
$29/11,3,23/11,2,17/11.$
We get the Hodge diamond
[…]
$\begin{array}[]{ccccccccccc}&&&&&1&&&&&\\ &&&&0&&0&&&&\\ &&&0&&1&&0&&&\\ &&0&&0&&0&&0&&\\ &0&&0&&2&&0&&0&\\ 0&&0&&4&&4&&0&&0\\ &0&&0&&2&&0&&0&\\ &&0&&0&&0&&0&&\\ &&&0&&1&&0&&&\\ &&&&0&&0&&&&\\ &&&&&1&&&&&\\ \end{array}+\begin{array}[]{ccccccccccc}&&&&&0&&&&&\\ &&&&0&&0&&&&\\ &&&0&&0&&0&&&\\ &&0&&0&&0&&0&&\\ &0&&1&&10&&1&&0&\\ 0&&0&&0&&0&&0&&0\\ &0&&0&&0&&0&&0&\\ &&0&&0&&0&&0&&\\ &&&0&&0&&0&&&\\ &&&&0&&0&&&&\\ &&&&&0&&&&&\\ \end{array}$
for the mixed Hodge structures $\operatorname{MHS}_{\operatorname{orb}}$ and $\operatorname{MHS}_{f}$.
Example 3.20
Let $f$ be the Laurent polynomial defined on $(\mathbb{C}^{*})^{4}$ by
$f(u)=u_{1}+u_{2}+u_{3}+u_{4}+\frac{1}{u_{1}^{5}u_{2}^{12}u_{3}^{12}u_{4}^{30}}.$
We have $\mu=60$. Its Newton polytope $P$ is the reduced and reflexive simplex of weight $(1,5,12,12,30)$ in $\mathbb{R}^{4}$. We get the Hodge diamond
$\begin{array}[]{ccccccccc}&&&&1&&&&\\ &&&0&&0&&&\\ &&0&&6&&0&&\\ &0&&5&&5&&0&\\ 0&&4&&18&&4&&0\\ &0&&5&&5&&0&\\ &&0&&6&&0&&\\ &&&0&&0&&&\\ &&&&1&&&&\end{array}.$
The Hodge diamond is symmetric about the horizontal axis because $P$ is reflexive, see Remark 3.16.
4 Interlude: the Hodge-Tate condition
In this section, we study the numerical condition appearing in Corollary 3.14. We first recall various characterizations of mixed Hodge structures of Hodge-Tate type.
Lemma 4.1
Let $(H,F^{\bullet},W_{\bullet})$ be a mixed Hodge structure such that $W_{\bullet}$ is the weight filtration of a nilpotent endomorphism $N$ of $H$ centered at $r$. The following are equivalent:
  1. 1.
    $W_{2i+1}H=W_{2i}H$ for $i\in\mathbb{Z}$ and $\operatorname{gr}_{F}^{p}\operatorname{gr}_{2q}^{W}H=0\ \mbox{for}\ p\neq q$ (the filtrations $F^{\bullet}H$ and $W_{2\bullet}$ are opposite).
  2. 2.
    the Hodge numbers of the mixed Hodge structure satisfy $h^{p,q}=0$ for $p\neq q$,
  3. 3.
    $H=\oplus_{k}F^{k}H\cap W_{2k}H$,
  4. 4.
    we have an isomorphism
    $N^{r-2k}:\operatorname{gr}_{F}^{r-k}H\longrightarrow\operatorname{gr}^{k}_{F}H$ (28)
    for any integer $k$ such that $0\leq k\leq[r/2]$.
When these equivalent conditions are satisfied, we will say that the mixed Hodge structure $(H,F^{\bullet},W_{\bullet})$ is of Hodge Tate type.
Proof.  The equivalence between the first and the third point can be found in [7, Proposition 2.1.5], and the equivalence between the first and the last point is precisely [30, Corollary 2.3]. The remaining assertion is straightforward.   $\Box$
Therefore, and after Corollary 3.14, we naturally offer the following:
Definition 4.2
We will call the condition
[…]
the Hodge-Tate condition.
When the orbifold cohomology is integrally graded, this Hodge-Tate condition already appears in [15]. It also appears in this form in [9], and one can refer to loc. cit. in order to have an idea of how common this condition is.
The Hodge-Tate condition deserves attention for various reasons. First, if it is satisfied, it follows from the proof of Theorem 3.10 that one can use the standard conjugation in order to get a mixed Hodge structure (we denote by $H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{R})$ the corresponding real form):
Proposition 4.3
Assume that $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ is equipped with the standard conjugation $c$ defined in Section 3.2. Assume moreover that the Hodge-Tate condition (29) is satisfied. Then, the tuple
$\operatorname{MHS}_{\operatorname{orb}}:=(H_{\operatorname{orb}}^{2*}(\mathcal{X}),H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{R}),F^{\bullet},W_{\bullet})$
is a mixed Hodge structure.
In the Hodge-Tate case, we have also an easy description of the weight filtration:
Proposition 4.4
Let $W_{\bullet}$ be the weight filtration of a mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ defined by Theorem 3.10 and assume that the Hodge-Tate condition (29) is satisfied. Then,
$W_{2k}H_{\operatorname{orb}}^{2*}(\mathcal{X})=W_{2k+1}H_{\operatorname{orb}}^{2*}(\mathcal{X})=\oplus_{n-k-1<\alpha}H^{2\alpha}_{\operatorname{orb}}(\mathcal{X})$ (30)
for $m\in\mathbb{Z}$.
Proof.  The proof of the first assertion is analogous to the one of Proposition 3.2: one uses (19), (20) and the fact that, under the assumption (29), $n_{\ell}=n-2i_{\ell}$ if $i_{\ell}\in\mathbb{Z}$ and $n_{\ell}=n-2[i_{\ell}]-1$ if $i_{\ell}\notin\mathbb{Z}$ in order to get
$W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})=\oplus_{2i\leq m}\oplus_{n-i-1<\alpha\leq n-i}H_{\operatorname{orb}}^{2\alpha}(\mathcal{X})$
for $m\in\mathbb{Z}$.   $\Box$
Corollary 4.5
Assume that $i_{\ell}\in\mathbb{Z}$ for all $\ell\in F$. Then,
$\operatorname{gr}^{W}_{2(n-k)}H_{\operatorname{orb}}^{2*}(\mathcal{X})=H_{\operatorname{orb}}^{2k}(\mathcal{X})$
if and only if the Hodge-Tate condition (29) is satisfied.
Proof.  The if part follows from Proposition 4.4. For the only if part, we use the description of the Hodge filtration given by Proposition 3.2 and Lemma 4.1.   $\Box$
Last, the Hodge-Tate condition is also connected with hard Lefschetz properties for orbifold cohomology. In what follows, we write $N$ for $L_{\omega}$, the nilpotent endomorphism of $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ defined by (12). we will say that $H_{\operatorname{orb}}^{2*}(\mathcal{X})$ satisfies the hard Lefschetz property if the powers of $N$ induces isomorphisms
$N^{n-2k}:H_{\operatorname{orb}}^{2k}(\mathcal{X})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H_{\operatorname{orb}}^{2(n-k)}(\mathcal{X})$ (31)
for all integers $k$ such that $0\leq k\leq n/2$ and
$N^{n-1-2k}:H^{2(\beta+k)}_{\operatorname{orb}}(\mathcal{X})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{2(\beta+n-1-k)}_{\operatorname{orb}}(\mathcal{X})$ (32)
for all $\beta\in]0,1[$ and integers $k$ such that $0\leq k\leq(n-1)/2$ (see also [30]). The next result (compare with [15, Proposition 4.1]), can be found with a different proof in [9, Theorem 3.4]. It is useful in order to test whether the orbifold cohomology satisfies a Kähler package, see Remark 5.12.
Proposition 4.6
The hard Lefschetz properties (31) and (32) hold if and only if the Hodge-Tate condition (29) is satisfied.
Proof.  By the description of the Hodge filtration given by Proposition 3.2, we have
$\operatorname{gr}^{n-1-k}_{F}H^{2*}_{\operatorname{orb},\neq 0}(\mathcal{X})=\bigoplus_{0<\beta<1}H^{2(\beta+k)}_{\operatorname{orb}}(\mathcal{X})\ \mbox{and}\ \operatorname{gr}^{k}_{F}H^{2*}_{\operatorname{orb},\neq 0}(\mathcal{X})=\bigoplus_{0<\beta<1}H^{2(\beta+n-1-k)}_{\operatorname{orb}}(\mathcal{X})$
therefore (32) holds if and only if
$N^{n-1-2k}:\operatorname{gr}^{n-1-k}_{F}H^{2*}_{\operatorname{orb},\neq 0}(\mathcal{X})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{gr}^{k}_{F}H^{2*}_{\operatorname{orb},\neq 0}(\mathcal{X})$
that is if and only if $i_{\ell}]=[i_{\ell^{-1}}]$ for all $\ell\in F$ such that $i_{\ell}\notin\mathbb{Z}$ by Corollary 3.14 and Lemma 4.1. In the same way, we get that (31) is true if and only if $i_{\ell}=i_{\ell^{-1}}$ for all $\ell\in F$ such that $i_{\ell}\in\mathbb{Z}$.   $\Box$
Remark 4.7
Assume that the Hodge-Tate condition (29) is satisfied and that $i_{\ell}\in\mathbb{Z}$ for all $\ell\in F$. Then,
$F^{p}H_{\operatorname{orb}}^{2*}(\mathcal{X})=\bigoplus_{\alpha\leq n-p}H_{\operatorname{orb}}^{2\alpha}(\mathcal{X})\ \mbox{and}\ W_{m}H_{\operatorname{orb}}^{2*}(\mathcal{X})=\bigoplus_{2i\leq m}H_{\operatorname{orb}}^{2(n-i)}(\mathcal{X})$
and, in our (simplicial) toric setting, our filtrations agree with the ones defined in [15, Section 5]. In this case, the mixed Hodge structure of Proposition 4.3 coincide with the one defined in [15, Theorem 5.3].
5 Polarizations
We now explain how the orbifold Poincaré duality defined by Chen and Ruan in [3] provides a polarization of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ defined in Theorem 3.10. We consider two special cases, leading to different statements: the first one, discussed in Section 5.1, looks like the duality of $\mathcal{D}$-modules introduced in [29] (see Appendix B.1) and fits with Sabbah’s mixed Hodge structures; the second one, discussed in Section 5.2 and defined only in the Hodge-Tate case, is more classical and follows along the lines the approach used in [15].
5.1 Orbifold Poincaré-Saito duality and polarization
In this section, we assume that $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is equipped with the Poincaré-Lefschetz conjugation (24). Using the orbifold Poincaré duality defined in [3], we construct a polarization of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ in the sense of Definition A.3. This construction is basically inspired by the description of the duality of $\mathcal{D}$-module for weighted projective spaces given in Appendix B.2. In what follows, we will set r=n1r=n-1 when considering $H^{2*}_{\operatorname{orb}}(\mathcal{X})_{\neq 0}$ and $r=n$ when considering $H^{2*}_{\operatorname{orb}}(\mathcal{X})_{0}$. We will denote by $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ either $H^{2*}_{\operatorname{orb}}(\mathcal{X})_{0}$ or $H^{2*}_{\operatorname{orb}}(\mathcal{X})_{\neq 0}$.
5.1.1 Orbifold Poincaré-Saito duality
After [3, Proposition 3.3.1], and for $0\leq\alpha\leq n$, we define the bilinear form
$g:H^{2\alpha}_{\operatorname{orb}}(\mathcal{X})\times H^{2(n-\alpha)}_{\operatorname{orb}}(\mathcal{X})\rightarrow\mathbb{C}$
as the direct sum over the sectors $X_{\ell}$ such that $\alpha-i_{\ell}$ is a nonnegative integer of
$g^{(\ell)}:H^{2(\alpha-i_{\ell})}(X_{\ell})\times H^{2(n-\alpha-i_{\ell^{-1}})}(X_{\ell^{-1}})\rightarrow\mathbb{C}$
where
$g^{(\ell)}(a,b)=\frac{(-1)^{[\alpha]}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge I^{*}(b),$ (33)
II denoting the involution $I:X_{\ell}\rightarrow X_{\ell^{-1}}$ alluded to in Section 2.2. We refer to [3] or [23, III 3.3] for the definition of the orbifold integral. We will call $g$ the Poincaré-Saito duality.
We first list the basic properties of the Poincaré-Saito duality, keeping in mind polarizations of mixed Hodge structures (see Appendix A).
Lemma 5.1
Let $g$ be the Poincaré-Saito duality and assume that $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is equipped with the Poincaré-Lefschetz conjugation (24). Then,
  1. 1.
    $g$ is nondegenerate,
  2. 2.
    $g$ is $(-1)^{r}$-symmetric,
  3. 3.
    $\overline{g(a,b)}=g(\overline{a},\overline{b})$ if $a\in H^{2(\alpha-i_{\ell})}(X_{\ell})$ and $b\in H^{2(n-\alpha-i_{\ell^{-1}})}(X_{\ell^{-1}})$,
  4. 4.
    $g(2i\pi Na,b)+g(a,2i\pi Nb)=0$.
Proof.  The proof basically relies on the normalization of the integral in the definition of $g^{(\ell)}$.
  1. 1.
    By (6), we have $n-\alpha-i_{\ell^{-1}}=n_{\ell}-(\alpha-i_{\ell})$ therefore $g^{(\ell)}$ is isomorphic to the ordinary Poincaré duality on $X_{\ell}$. In particular, $g$ is nondegenerate since $g^{(\ell)}$ is.
  2. 2.
    Let $a\in H^{2(\alpha-i_{\ell})}(X_{\ell})$ and let $b\in H^{2(n-\alpha-i_{\ell^{-1}})}(X_{\ell^{-1}})$. Then,
    $g(b,a)=\frac{(-1)^{[n-\alpha]}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell^{-1}}}^{\operatorname{orb}}b\wedge I^{*}(a)=\frac{(-1)^{r-[\alpha]}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}I^{*}(b)\wedge a=(-1)^{r}g(a,b).$
  3. 3.
    Let $a\in H^{2(\alpha-i_{\ell})}(X_{\ell})$ and let $b\in H^{2(n-\alpha-i_{\ell^{-1}})}(X_{\ell^{-1}})$. Then,
    $\overline{g(a,b)}=(-1)^{r}\frac{(-1)^{[\alpha]}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}c(a)\wedge I^{*}(c(b))$
    $\frac{(-1)^{[\alpha]-[i_{\ell^{-1}}]-[i_{\ell}]}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell^{-1}}}^{\operatorname{orb}}\overline{a}\wedge I^{*}(\overline{b})=\frac{(-1)^{[\beta]}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell^{-1}}}^{\operatorname{orb}}\overline{a}\wedge I^{*}(\overline{b})=g(\overline{a},\overline{b})$
    because $I^{*}(\overline{a})=(-1)^{\alpha-i_{\ell}}c(a)$, $I^{*}(\overline{b})=(-1)^{n-\alpha-i_{\ell^{-1}}}c(b)$, $n-i_{\ell^{-1}}-i_{\ell}=r-[i_{\ell^{-1}}]-[i_{\ell}]$ and $\overline{a}\in H^{2(\beta-i_{\ell^{-1}})}(X_{\ell^{-1}})$ with $\beta-i_{\ell^{-1}}=\alpha-i_{\ell}\in\mathbb{Z}$.
  4. 4.
    Let $a\in H^{2(\alpha-i_{\ell})}(X_{\ell})$. Then,
    $g(2i\pi Na,b)=\frac{(-1)^{[\alpha]+1}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}2i\pi N(a)\wedge I^{*}(b)=\frac{(-1)^{[\alpha]+1}}{(2i\pi)^{r}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge I^{*}(2i\pi N(b))$
    $g(a,2i\pi Nb)$
    where the second equality follows from Lemma 3.9.
This completes the proof.   $\Box$
The following result provides useful orthogonality relations:
Lemma 5.2
Let $F^{\bullet}$ and $W_{\bullet}$ be the Hodge filtration of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$ and let $g$ be the Poincaré-Saito duality. Then,
  1. 1.
    $g(F^{p},F^{r+1-p})=0$,
  2. 2.
    $g(W_{i},W_{2r-i-1})=0$.
Proof.  The first point follows from Proposition 3.2 and from the definition of $g$. The second point about the weight filtration can be directly shown as follows: we have
$W_{i}=\oplus_{r+n_{\ell}-i\leq 2j}H^{2j}(X_{\ell})\ \mbox{and}\ W_{2r-i-1}=\oplus_{n_{\ell}+i+1-r\leq 2j^{\prime}}H^{2j^{\prime}}(X_{\ell})$
therefore, if $j$ (resp. $j^{\prime}$) is an index appearing in the first (resp. second) direct sum decomposition, we have $2j+2j^{\prime}\geq 2n_{\ell}+1$. The assertion follows.   $\Box$
5.1.2 A polarization
For a nonnegative integer $\nu$, we define the primitive subspaces
$P_{r+\nu}:=\ker((2i\pi N)^{\nu+1}:\operatorname{gr}^{W}_{r+\nu}H^{2*}_{\operatorname{orb}}(\mathcal{X})\rightarrow\operatorname{gr}^{W}_{r-\nu-2}H^{2*}_{\operatorname{orb}}(\mathcal{X})).$
We now show that the Poincaré-Saito duality $g$ provides a polarization of the induced Hodge structure
$P_{r+\nu}=\oplus_{p}F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}}.$ (34)
First, and thanks to Lemma 5.1, it is possible to apply the construction recalled in the appendix: we define the nondegenerate bilinear form $g_{\nu}$ on $\operatorname{gr}^{W}_{r+\nu}H^{2*}_{\operatorname{orb}}(\mathcal{X})$ by putting
[…]
if $\tilde{a},\tilde{b}\in W_{r+\nu}H^{2*}_{\operatorname{orb}}(\mathcal{X})$ represent $a,b\in\operatorname{gr}^{W}_{r+\nu}H^{2*}_{\operatorname{orb}}(\mathcal{X})$ and the sesquilinear form
$h_{\nu}:P_{r+\nu}\times P_{r+\nu}\rightarrow\mathbb{C}$
by putting
$h_{\nu}(u,v):=i^{2p-(r+\nu)}g_{\nu}(u,\overline{v})$ (35)
if $u\in F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}}$.
Lemma 5.3
Let $\nu$ be a nonnegative integer.
  1. 1.
    The Hodge subspaces of the Hodge decomposition (34) are orthogonal with respect to the sesquilinear form $h_{\nu}$.
  2. 2.
    The sesquilinear form $h_{\nu}$ is Hermitian.
Proof.  By Lemma 5.2, we have $g_{\nu}(F^{p}P_{r+\nu},F^{r+\nu+1-p}P_{r+\nu})=0$ and, by Lemma 5.1, these orthogonality relations are preserved under conjugation. Therefore
$h_{\nu}(F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}},F^{q}P_{r+\nu}\cap\overline{F^{r+\nu-q}P_{r+\nu}})=0$
if $r+\nu-q\geq r+\nu+1-p$ or if $q\geq r+\nu+1-(r+\nu)+p$ that is if $q\leq p-1$ or $q\geq p+1$. For the second point, we use the previous assertion, the reality of $g$ and the $(-1)^{r+\nu}$-symmetry of $g_{\nu}$.   $\Box$
Lemma 5.4
Let $\nu$ be a nonnegative integer.
  1. 1.
    We have $P_{r+\nu}=\bigoplus_{\ell\in F}H^{n_{\ell}-\nu}_{\operatorname{prim}}(X_{\ell}):=\bigoplus_{\ell\in F}\ker(N^{\nu+1}:H^{n_{\ell}-\nu}(X_{\ell})\rightarrow H^{n_{\ell}+\nu+2}(X_{\ell}))$.
  2. 2.
    Let $a\in H^{n_{\ell}-\nu}_{\operatorname{prim}}(X_{\ell})$. Then, $a\in F^{r-[i_{\ell}]-k}P_{r+\nu}\cap\overline{F^{r+\nu-(r-[i_{\ell}]-k)}P_{r+\nu}}$ where $2k=n_{\ell}-\nu$.
  3. 3.
    Let $a\in H^{n_{\ell}-\nu}_{\operatorname{prim}}(X_{\ell})$. Then,
    $h_{\nu}(a,a)=\frac{1}{(2\pi)^{r-\nu}}(-1)^{\alpha-i_{\ell}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge N^{\nu}c(a)$ (36)
    where $2(\alpha-i_{\ell})=2k=n_{\ell}-\nu$.
In particular, $h_{\nu}(a,a)>0$ if $a\in F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}}$ and $a\neq 0$.
Proof.  We keep the notations of the statement of the lemma.
  1. 1.
    By the very definition of the weight filtration (see Section 3.1.2), we have
    $\operatorname{gr}^{W}_{r+\nu}H_{\operatorname{orb}}^{2*}=\bigoplus_{\ell\in F}H^{n_{\ell}-\nu}(X_{\ell})$
    whence the first assertion, because NN preserves each sector (see Remark 2.6).
  2. 2.
    By the very definition of the Hodge filtration (see Section 3.1.1), we have $a\in F^{r-[i_{\ell}]-k}P_{r+\nu}$. Using (24), we get $\overline{a}\in F^{r-[i_{\ell^{-1}}]-k}P_{r+\nu}$ therefore
    $a\in\overline{F^{r-[i_{\ell^{-1}}]-k}P_{r+\nu}}=\overline{F^{r+\nu-(r-[i_{\ell}]-k)}P_{r+\nu}}$
    because $r-[i_{\ell^{-1}}]-k=\nu+[i_{\ell}]+k$ by (6).
  3. 3.
    From the second point and the definition of $h_{\nu}$, we get (keeping in mind the definition of the Poincaré-Lefschetz conjugation (24))
    $h_{\nu}(a,a)=i^{[i_{\ell^{-1}}]-[i_{\ell}]}g(a,(2i\pi N)^{\nu}\overline{a})=(2i\pi)^{\nu}i^{[i_{\ell^{-1}}]-[i_{\ell}]}g(a,N^{\nu}\overline{a})$
    $(2i\pi)^{\nu-r}i^{[i_{\ell^{-1}}]-[i_{\ell}]}(-1)^{[\alpha]}(-1)^{\alpha-i_{\ell}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge N^{\nu}c(a)=\frac{1}{(2\pi)^{r-\nu}}(-1)^{\alpha-i_{\ell}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge N^{\nu}c(a)$
    because $r-\nu=[i_{\ell}]+[i_{\ell^{-1}}]+2(\alpha-i_{\ell})$ thanks again to (6).
Now, the right hand term of (36) is positive if $a\neq 0$ (see for instance [15, Theorem 2.2], [33], [38]; the standard Hodge structure on $X_{\ell}$ is of Hodge-Tate type). The last assertion follows since, by the second point, $F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}}$ is a suitable direct sum of primitive vector spaces $H^{2k}_{\operatorname{prim}}(X_{\ell})$ and $h_{\nu}$ preserves this direct sum decomposition.   $\Box$
Theorem 5.5
Assume that $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is equipped with the Poincaré-Lefschetz conjugation (24) and let $g$ be the Poincaré-Saito orbifold duality. Then, the tuple
$(H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{C}),H_{\operatorname{orb},\mathbb{R}}^{2*}(\mathcal{X}),F^{\bullet},W_{\bullet},N,g,r)$
is a polarized mixed Hodge structure of weight $r$ in the sense of Definition A.3 (see the appendix).
Proof.  Properties 1-5 of Definition A.3 are now clear (Lemma 5.1 provides the required properties for $g$). Property 6 follows from Lemma 5.3 and property 7 follows from Lemma 5.4.   $\Box$
It is interesting to specialize the previous results to weighted projective spaces, for which we do not need to refer to any conceptual result (in particular the hard Lefschetz property) in order to get a polarization.
Proposition 5.6
Let $P$ be a reduced simplex of weight $q=(q_{0},\ldots,q_{n})$ as in Example 2.8 and let $\nu$ be a nonnegative integer.
  1. 1.
    We have $P_{r+\nu}=\bigoplus_{\ell\in F,\ n_{\ell}=\nu}H^{0}(X_{\ell})$.
  2. 2.
    Let $a\in H^{0}(X_{\ell})$ with $n_{\ell}=\nu$. Then $a\in F^{r-[i_{\ell}]}P_{r+\nu}\cap\overline{F^{r+\nu-(r-[i_{\ell}])}P_{r+\nu}}$.
  3. 3.
    Let $a\in H^{0}(X_{\ell})$ with $n_{\ell}=\nu$. Then,
    $h_{\nu}(a,a)=\frac{1}{(2\pi)^{r-\nu}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge N^{\nu}c(a).$ (37)
In particular, the Poincaré-Saito duality provides a polarization of the mixed Hodge structure $\operatorname{MHS}_{\operatorname{orb}}$.
Proof.  Recall that $\operatorname{gr}^{W}_{r+\nu}H_{\operatorname{orb}}^{2*}=\oplus_{\ell\in F}H^{n_{\ell}-\nu}(X_{\ell})$. By [21] we have $\dim H^{2i}(X_{\ell})=1$ and $H^{2i+1}(X_{\ell})=0$ for all $\ell\in F$ therefore the map
$N^{\nu+1}:H^{n_{\ell}-\nu}(X_{\ell})\rightarrow H^{n_{\ell}+\nu+2}(X_{\ell})$
is injective if $\nu\leq n_{\ell}-2$ since $N^{n_{\ell}}\neq 0$ (see the remark below). This shows the first assertion. The remaining ones are shown as in Lemma 5.4.   $\Box$
Remark 5.7
In order to calculate the integrals (37), we use [23]: for instance, with the notations of loc. cit. we have $\omega:=\varphi^{-1}([f])=\mu c_{1}^{\operatorname{orb}}(O_{\mathbb{P}(q)}(1))$ and
$\int^{\operatorname{orb}}_{\mathbb{P}(q)}N^{n}=\mu^{n}\int_{\mathbb{P}(q)}^{\operatorname{orb}}(c_{1}^{\operatorname{orb}}(O_{\mathbb{P}(q)}(1)))^{n}=\frac{\mu^{n}}{q_{0}q_{1}\ldots q_{n}}$
by [23, Proposition IV.3.13] (this is the volume of the polar polytope $P^{*}$ of $P$; recall that $\mu=q_{0}+q_{1}+\ldots+q_{n}$). Since the twisted sectors $\mathcal{X}_{\ell}$ are also weighted projective spaces, we have analogous formulas for $\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}N^{n_{\ell}}$: for $\ell\in F$ we have
$\int^{\operatorname{orb}}_{\mathcal{X}_{\ell}}N^{n_{\ell}}=\frac{\mu^{n_{\ell}}}{\prod_{i\in I(\ell)}q_{i}}$
where $I(\ell):=\{i,\ q_{i}\ell\in\mathbb{Z}\}$. On the way, we get the hard Lefschetz isomorphisms
$N^{n_{\ell}-2k}:H^{2k}(X_{\ell})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{2(n_{\ell}-k)}(X_{\ell})$
for $2k\leq n_{\ell}$.
Corollary 5.8
Let $P$ be a reduced simplex and let $\nu$ be a nonnegative integer. Then,
$P_{r+\nu}=\bigoplus_{p\in\mathbb{Z}}F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}}$
where
$F^{p}P_{r+\nu}\cap\overline{F^{r+\nu-p}P_{r+\nu}}=\bigoplus_{\ell\in F,\ n_{\ell}=\nu,\ [i_{\ell}]=r-p}H^{0}(X_{\ell}).$
5.2 Alternate outtake: orbifold Poincaré-Fernandez duality
In this section, we assume that $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is equipped with the Poincaré-Lefschetz conjugation (25): in this case, and by Lemma 3.9, $N$ is real while $2i\pi N$ is not, so we leave the singularity framework, but this setting may be of interest since it very close from the one considered by Fernandez in [15] and goes along with the existence of Kähler packages (see Remark 5.12) on the orbifold cohomology. However, it should be emphasized that in this case we get a polarization only under the Hodge-Tate condition (29).
For $0\leq\alpha\leq n$, we define the bilinear forms
$g_{\operatorname{PF}}:H^{2\alpha}_{\operatorname{orb}}(\mathcal{X})\times H^{2(n-\alpha)}_{\operatorname{orb}}(\mathcal{X})\rightarrow\mathbb{C}$
as the direct sum over the sectors $X_{\ell}$ such that $\alpha-i_{\ell}$ is a nonnegative integer of
$g^{(\ell)}_{\operatorname{PF}}:H^{2(\alpha-i_{\ell})}(X_{\ell})\times H^{2(n-\alpha-i_{\ell^{-1}})}(X_{\ell^{-1}})\rightarrow\mathbb{C}$
where now
$g^{(\ell)}_{\operatorname{PF}}(a,b)=(-1)^{\alpha-i_{\ell}}\int_{\mathcal{X}_{\ell}}^{\operatorname{orb}}a\wedge I^{*}(b).$ (38)
We will call it the Poincaré-Fernandez duality. We have the counterpart of Lemma 5.1 (its proof is analogous): the main difference is that $g_{\operatorname{PF}}$ is not $(-1)^{r}$-symmetric in general.
Lemma 5.9
The Poincaré-Fernandez duality satisfies the following properties:
  1. 1.
    $g_{\operatorname{PF}}$ is nondegenerate,
  2. 2.
    $g_{\operatorname{PF}}(b,a)=(-1)^{r-[i_{\ell^{-1}}]-[i_{\ell}]}g_{\operatorname{PF}}(a,b)$ if $a\in H^{2(\alpha-i_{\ell})}(X_{\ell})$ and $b\in H^{2(n-\alpha-i_{\ell^{-1}})}(X_{\ell^{-1}})$,
  3. 3.
    $g_{\operatorname{PF}}$ is real with respect to the Poincaré-Lefschetz conjugation (25),
  4. 4.
    $g_{\operatorname{PF}}(Na,b)+g_{\operatorname{PF}}(a,Nb)=0$.   $\Box$
We define the sequilinear form $h_{\nu}^{\operatorname{PF}}$ on $P_{r+\nu}$ as in Section 5.1, replacing $2i\pi N$ by $N$. Lemma 5.2 and Lemma 5.3 are still valid for $g_{\operatorname{PF}}$.
Lemma 5.10
Assume that the Hodge-Tate condition (29) holds. Then, the sesquilinear form $h_{\nu}^{\operatorname{PF}}$ is Hermitian and positive definite on the primitive space $P_{r+\nu}$.
Proof.  Under the Hodge-Tate type assumption, and by Lemma 5.9, the bilinear form $g_{\operatorname{PF}}$ is $(-1)^{r}$-symmetric therefore $h_{\nu}^{\operatorname{PF}}$ is Hermitian as in Lemma 5.3 and the remaining positivity assertion is shown as in Theorem 5.5.   $\Box$
Theorem 5.11
Assume that $H^{2*}_{\operatorname{orb}}(\mathcal{X})$ is equipped with a Poincaré-Lefschetz conjugation (25) and that the Hodge-Tate condition (29) holds. Then, the tuple
$(H_{\operatorname{orb}}^{2*}(\mathcal{X},\mathbb{C}),H_{\operatorname{orb},\mathbb{R}}^{2*}(\mathcal{X}),F^{\bullet},W_{\bullet},N,g_{\operatorname{PF}},r)$
is a polarized mixed Hodge structure of weight $r$ in the sense of Definition A.3.   $\Box$
Proof.  The statement follows from Lemma 5.9 and Lemma 5.10.   $\Box$
Remark 5.12
The previous construction is connected with the existence of a Kähler package in the sense of [1] for instance. Indeed, let us assume that the graded Milnor ring $A^{*}$ is integrally graded. Using the results of Section 4 and Section 5.2, it is possible to show that it satisfies a Kähler package in the sense of [1, Section 7] (with respect to $[f]$, the class of $f$ in $A^{1}$) if and only if the Hodge-Tate condition (29) holds (note that only the Hodge-Tate case is considered in loc. cit.). The corresponding Hodge-Riemann property probably provides the log-concavity of some relevant sequences (see [1], [18], [36], [37]…).
Appendix A Appendix (polarized mixed-Hodge structures)
We recall the definition of polarized (mixed-)Hodge structures used in these notes.
Definition A.1
A polarized Hodge structure of weight $r$ is a tuple $(H,H_{\mathbb{R}},F^{\bullet},g,r)$ where
  • $r$ is an integer,
  • $H_{\mathbb{R}}$ is a finite dimensional $\mathbb{R}$-vector space,
  • $H=H_{\mathbb{R}}\otimes\mathbb{C}$,
  • $g$ is a nondegenerate, $(-1)^{r}$-symmetric bilinear form on $H$, with real values on $H_{\mathbb{R}}$,
  • $F^{\bullet}$ is a decreasing and exhaustive filtration of $H$ (the Hodge filtration)
such that
$H=\oplus F^{p}\cap\overline{F^{r-p}},$ (39)
$g(F^{p}H,F^{r+1-p}H)=0,$ (40)
$i^{2p-r}g(u,\overline{u})>0\ \mbox{for all}\ u\in F^{p}H\cap\overline{F^{r-p}H},\ u\neq 0.$ (41)
We will also say that $g$ is a polarization of the pure Hodge structure $(H,H_{\mathbb{R}},F^{\bullet})$ of weight $r$.
Remark A.2
Let $h$ be the Hermitian form on $H$ defined by
$h(u,v):=i^{2p-r}g(u,\overline{v})$
if $u\in H^{p,r-p}:=F^{p}H\cap\overline{F^{r-p}H}$. By (40) the Hodge subspaces $H^{p,r-p}$ are orthogonal with respect to $h$ (first Hodge-Riemann bilinear relations) and by (41) $h$ is positive definite (second Hodge-Riemann bilinear relations).
Let $N:V\rightarrow V$ be a linear map on a finite dimensional vector space over a field of characteristic 0, which satisfies $N^{r+1}=0$. Then (see [35, Lemma 6.4]),
  • there exists a unique increasing filtration $W_{\bullet}$ of $V$
    $0\subset W_{0}\subset W_{1}\subset\cdots\subset W_{2r}=V$
    such that $N(W_{i})\subset W_{i-2}$ and
    $N^{\ell}:W_{r+\ell}/W_{r+\ell-1}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}W_{r-\ell}/W_{r-\ell-1}$ (42)
    This is the weight filtration of $N$ centered at $r$.
  • We have the Lefschetz decomposition
    $\operatorname{gr}_{r+\ell}^{W}V=\oplus_{i\geq 0}N^{i}P_{r+\ell+2i}$ (43)
    if $\ell\geq 0$ where the (primitive) subspaces $P_{r+\ell}$ of $\operatorname{gr}^{W}_{r+\ell}V$ are defined by
    $P_{r+\ell}:=\ker(N^{\ell+1}:\operatorname{gr}^{W}_{r+\ell}V\rightarrow\operatorname{gr}^{W}_{r-\ell-2}V)$ (44)
    if $\ell\geq 0$ and $P_{r+\ell}=0$ if $\ell<0$.
  • If $N$ is an infinitesimal isometry of a $(-1)^{r}$-symmetric nondegenerate linear form $g$ on $V$, i.e
    $g(Na,b)+g(a,Nb)=0$ (45)
    for all $a,b\in V$, then $g(W_{i},W_{i^{\prime}})=0$ for $i+i^{\prime}<2r$ and one can define a nondegenerate $(-1)^{r+\ell}$-symmetric bilinear form $g_{\ell}$ on $\operatorname{gr}^{W}_{r+\ell}V$ for $\ell\geq 0$ by
[…]
    where $\tilde{a}\in W_{r+\ell}$ (resp. $\tilde{b}\in W_{r+\ell}$) is a representative of $a$ (resp. $b$). By (45), the Lefschetz decomposition (43) is orthogonal with respect to $g_{\ell}$.
We apply the previous construction to the following situation: $V:=H_{\mathbb{R}}$ is a finite dimensional $\mathbb{R}$-vector space and $H$ is the complexification of $H_{\mathbb{R}}$.
Definition A.3
A polarized mixed Hodge structure of weight $r$ is a tuple $(H,H_{\mathbb{R}},F^{\bullet},g,r,W_{\bullet},N)$ such that:
  1. 1.
    $g$ is $(-1)^{r}$-symmetric nondegenerate bilinear form on $H$ with real values on $H_{\mathbb{R}}$,
  2. 2.
    the nilpotent endomorphism $N$ of $H_{\mathbb{R}}$ is an infinitesimal isometry of $g$,
  3. 3.
    $W_{\bullet}$ is the weight filtration of $N$ centered at $r$,
  4. 4.
    the filtration $F^{\bullet}\operatorname{gr}^{W}_{m}H$ defines a pure Hodge structure of weight $m$ on $\operatorname{gr}^{W}_{m}H$,
  5. 5.
    $N(F^{p})\subset F^{p-1}$,
  6. 6.
    $g(F^{p}H,F^{r+1-p}H)=0$,
  7. 7.
    $i^{2p-(r+\ell)}g_{\ell}(u,\overline{u})>0$ for all $u\in F^{p}P_{r+\ell}\cap\overline{F^{r+\ell-p}P_{r+\ell}}$, $u\neq 0$.
We will say that the tuple is pre-polarized if only the first six properties are satisfied.
Remark A.4
Let $(H,H_{\mathbb{R}},F^{\bullet},g,r,W_{\bullet},N)$ be a polarized mixed Hodge structure of weight $r$. The primitive subspace $P_{r+\ell}$ is a Hodge substructure as the kernel of a morphism of pure Hodge structures and we get the (induced) Hodge decomposition
$P_{r+\ell}=\oplus_{p}F^{p}P_{r+\ell}\cap\overline{F^{r+\ell-p}P_{r+\ell}}.$ (47)
This Hodge structure is polarized by $g_{\ell}$ in the sense of Remark A.2.
Appendix B Appendix: duality of $\mathcal{D}$-modules for weighted projective spaces
In this appendix, we explain why the construction of the polarization in Section 5.1 is inspired by the duality of $\mathcal{D}$-modules studied in [29] and applied in [14] to (mirror partners of) weighted projective spaces.
B.1 Foreword: about the duality of $\mathcal{D}$-modules
Given a $\mathbb{C}[\tau,\tau^{-1}]$-module $G$, we denote by $\overline{G}$ the $\mathbb{C}$-vector space $G$ equipped with the new module structure $p(\tau).g:=p(-\tau)g$. We will denote by $\overline{g}$ the elements of $\overline{G}$. If moreover $G$ is equipped with a connection $\partial_{\tau}$ then so is $G$ and we have $\partial_{\tau}\overline{g}:=\overline{-\partial_{\tau}g}$. When $G$ is the Fourier-transform of the Gauss-Manin system of $f$ as in Section 2.3, duality of $\mathcal{D}$-modules gives (see [29, Section 3] for details and also [32, Section 2.7] for the local model) the existence of nondegenerate $\mathbb{C}[\tau,\tau^{-1}]$-bilinear pairings
$S:G\times\overline{G}\longrightarrow\mathbb{C}[\tau,\tau^{-1}]$
such that
  1. 1.
    $\tau\partial_{\tau}S(g^{\prime},\overline{g^{\prime\prime}})=S(\tau\partial_{\tau}g^{\prime},\overline{g^{\prime\prime}})+S(g^{\prime},\overline{\tau\partial_{\tau}g^{\prime\prime}})$,
  2. 2.
    $S$ sends $G_{0}\times\overline{G_{0}}$ in $\mathbb{C}[\theta]\theta^{n}$,
  3. 3.
    $S(g^{\prime\prime},\overline{g^{\prime}})=(-1)^{n}\overline{S(g^{\prime},\overline{g^{\prime\prime}})}$.
Such a pairing $S$ induces nondegenerate bilinear forms
$g_{\mathcal{D}}:H_{\alpha}\times H_{1-\alpha}\longrightarrow\mathbb{C}\tau^{-1}\ \mbox{for}\ \alpha\neq 0\ \mbox{and}\ g_{\mathcal{D}}:H_{0}\times H_{0}\longrightarrow\mathbb{C}.$
This provides a $(-1)^{n-1}$-symmetric nondegenerate bilinear form $g_{\mathcal{D}}$ on $H_{\neq 0}$ and a $(-1)^{n}$-symmetric nondegenerate bilinear form on $H_{0}$ such that
$g_{\mathcal{D}}(F^{p},F^{r+1-p})=0,$
where $F^{\bullet}$ denotes the Hodge filtration of the mixed Hodge structure $\operatorname{MHS}_{f}$ (as always, $r=n$ on $H_{0}$ and r=n1r=n-1 otherwise).
B.2 Application to weighted projective spaces
Let $P$ be a reduced simplex of weight $q=(q_{0},\ldots,q_{n})$ in the sense of Example 2.8 and let $f$ be the Laurent polynomial defined by
$f(u)=\sum_{b\in\mathcal{V}(P)}u^{b}$
as in (8). We first describe the pairing $g_{\mathcal{D}}$ on $H$. We follow closely [14].
Recall the classes $\omega_{0},\ldots,\omega_{\mu-1}$ in $G_{0}$ defined in [14, Proposition 3.2], which give a Birkhoff normal form for the Brieskorn lattice $G_{0}$. We will denote by $\alpha_{i}$ the $V$-order of $\omega_{i}$. By [14, Lemma 4.1], there exists a unique (up to a nonzero constant) pairing $S$ as above: it is given by the formulas (the coefficients of $f$ differs slightly from the one in loc. cit., and this explains the mild modifications in the formulas defining $S$):
$S(\omega_{k},\overline{\omega_{\ell}})=\left\{\begin{array}[]{ll}c_{k,\ell}S(\omega_{0},\overline{\omega_{n}})&\mbox{if}\ 0\leq k\leq n\ \mbox{and}\ k+\ell=n\\ c_{k,\ell}\frac{1}{q^{q}}S(\omega_{0},\overline{\omega_{n}})&\mbox{if}\ n+1\leq k\leq\mu-1\ \mbox{and}\ k+\ell=\mu+n\\ 0&\mbox{otherwise}\end{array}\right$. (48)
where $q^{q}=q_{0}^{q_{0}}\ldots q_{n}^{q_{n}}$ and $c_{k,\ell}$ is a positive rational number expressed with the help of the weights $q_{0},\ldots,q_{n}$, see for instance [23, Lemme VI.1.18]. Moreover, $S(\omega_{k},\overline{\omega_{\ell}})\in\mathbb{C}\tau^{-n}$ and we will write
$S(\omega_{0},\overline{\omega_{n}}):=S^{o}(\omega_{0},\omega_{n})\tau^{-n}.$
The pairing $S$ induces a nondegenerate bilinear form $g_{\mathcal{D}}$ on $H$ as follows (the brackets denote the class in $H$):
  • on $H_{\neq 0}$,
    $g_{\mathcal{D}}([\tau^{[\alpha_{i}]}\omega_{i}],[\tau^{[\alpha_{j}]}\omega_{j}])=(-1)^{[\alpha_{j}]}S(\omega_{i},\overline{\omega_{j}})\tau^{n-1}$ (49)
    if $\alpha_{i}]+[\alpha_{j}]=n-1$ and 0 otherwise,
  • on $H_{0}$,
    $g_{\mathcal{D}}([\tau^{\alpha_{i}}\omega_{i}],[\tau^{\alpha_{j}}\omega_{j}])=(-1)^{\alpha_{j}}S(\omega_{i},\overline{\omega_{j}})\tau^{n}$ (50)
    if $\alpha_{i}+\alpha_{j}=n$ and 0 otherwise.
This bilinear form is $(-1)^{n-1}$-symmetric on $H_{\neq 0}$, $(-1)^{n}$-symmetric on $H_{0}$.
Let $k$ be such that $\omega_{k}^{\prime}]:=[\tau^{[\alpha_{k}]}\omega_{k}]$ is a primitive element of $N$ (we will denote by $K$ the set of these indices): the space
$B_{k}:=<N^{j}[\omega_{k}^{\prime}],\ j=0,\ldots,\nu_{k}>$
is a Jordan block of $N$ of size $\nu_{k}+1$ and we will call the integer $\nu_{k}$ is the weight of $\omega_{k}^{\prime}]$. For $k\in K$ such that $k\geq n+1$, the class $\omega_{\mu+n-k-\nu_{k}}^{\prime}]$ is also a primitive vector of $N$ of weight $\nu_{k}$: we will put $\overline{k}:=\mu+n-k-\nu_{k}$ and $\overline{k+j}:=\overline{k}+j$ for $j=0,\ldots,\nu_{k}$ (if $k\in[0,n]$ we simply define $\overline{k}:=k$). Note that
$\alpha_{k}+\alpha_{\overline{k}}+\nu_{k}=n$ (51)
because the $V$-order of $\omega_{\mu+n-k}$ is equal to $n-\alpha_{k}$ and also to $\alpha_{\overline{k}}+\nu_{k}$ (compare with (6)). The next result is a motivation for the definition of Poincaré-Lefschetz conjugations (see Section 3.2).
Lemma B.1
[14, Lemma 5.3] Let $k\in K$. Then,
$\overline{[\omega_{k}^{\prime}]}=\sum_{\ell=k}^{k+\nu_{k}}a_{k,\overline{\ell}}[\omega_{\overline{\ell}}^{\prime}]$ (52)
where $\overline{a_{k,\overline{k}}}a_{\overline{k},k}=1$ (in particular, $a_{k,\overline{k}}\neq 0$).   $\Box$
In what follows, the double brackets denote the class in $\operatorname{gr}^{W}_{\bullet}H$ and we consider the bilinear form $g_{\mathcal{D},\ell}$ induced on $\operatorname{gr}^{W}_{r+\ell}$ by $g_{\mathcal{D}}$ as in (46), using the nilpotent endomorphism $2i\pi N$. The two next results must be compared with Lemma 5.4 and Theorem 5.5 respectively.
Proposition B.2
Let $\omega_{k}^{\prime}]$ be a primitive element of $N$ of weight $\nu_{k_{0}}$. Then,
  1. 1.
    $\omega_{k}^{\prime}]]\in P_{r+\nu_{k_{0}}}$ and $P_{r+\nu_{k_{0}}}$ is generated by such classes,
  2. 2.
    $\omega_{k}^{\prime}]]\in F^{r-[\alpha_{k}]}P_{r+\nu_{k_{0}}}\cap\overline{F^{r+\nu_{k_{0}}-(r-[\alpha_{k}])}P_{r+\nu_{k_{0}}}}$,
  3. 3.
    for any primitive element $\omega_{\ell}^{\prime}]$ of weight $\nu_{k_{0}}$,
    $g_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],\overline{[[\omega_{\ell}^{\prime}]]})=\left\{\begin{array}[]{ll}0&\mbox{if}\ \ell\neq k\\ a_{k,\overline{k}}(2i\pi)^{\nu_{k_{0}}}(-1)^{r-[\alpha_{k}]}c_{k,\mu+n-k}\frac{1}{q^{q}}S^{o}(\omega_{0},\omega_{n})&\mbox{if}\ \ell=k\neq 0\\ a_{0,0}(2i\pi)^{n}(-1)^{n}S^{o}(\omega_{0},\omega_{n})&\mbox{if}\ \ell=k=0\end{array}\right$. (53)
  4. 4.
    the Hodge subspaces $F^{p}P_{r+\nu_{k_{0}}}\cap\overline{F^{r+\nu_{k_{0}}-p}P_{r+\nu_{k_{0}}}}$ are orthogonal with respect to $g_{\mathcal{D},\nu_{k_{0}}}(\bullet,\overline{\bullet})$.
Proof.  The first point follows from the definition of the primitive subspaces. By the very definition of the Hodge filtration, we have $\omega_{k}^{\prime}]]\in F^{r-[\alpha_{k}]}P_{r+\nu_{k_{0}}}$. Now, we have also $\omega_{\overline{k}}^{\prime}]]\in P_{r+\nu_{k_{0}}}$ since $\nu_{k_{0}}=\nu_{\overline{k}_{0}}$ and it follows from (52) that $\overline{[[\omega_{k}^{\prime}]]}=a_{k,\overline{k}}[[\omega_{\overline{k}}^{\prime}]]\in F^{r-[\alpha_{\overline{k}}]}P_{r+\nu_{k_{0}}}$. Therefore,
$\omega_{k}^{\prime}]]\in\overline{F^{r-[\alpha_{\overline{k}}]}P_{r+\nu_{k_{0}}}}=\overline{F^{r+\nu_{k_{0}}-(r-[\alpha_{k}])}P_{r+\nu_{k_{0}}}}$
where the equality follows from (51). Formulas (53) follow from the equalities
$g_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],\overline{[[\omega_{\ell}^{\prime}]]})=a_{\ell,\overline{\ell}}(2i\pi)^{\nu_{k_{0}}}g_{\mathcal{D}}([\omega_{k}^{\prime}],N^{\nu_{k_{0}}}[\omega_{\overline{\ell}}^{\prime}])=a_{\ell,\overline{\ell}}(2i\pi)^{\nu_{k_{0}}}g_{\mathcal{D}}([\omega_{k}^{\prime}],[\omega_{\overline{\ell}+\nu_{k_{0}}}^{\prime}]),$
(48), (49) and (50). The fourth point follows from 1, 2 and 3.   $\Box$
Let $h_{\mathcal{D},\ell}$ be the sesquilinear form defined on the primitive subspace $P_{r+\ell}$ by
$h_{\mathcal{D},\ell}(u,v):=i^{2p-(r+\ell)}g_{\mathcal{D},\ell}(u,\overline{v})$ (54)
if $u\in F^{p}P_{r+\ell}\cap\overline{F^{r+\ell-p}P_{r+\ell}}$.
Proposition B.3
Let $\omega_{k}^{\prime}]$ be a primitive element of $N$ of weight $\nu_{k_{0}}$. Then,
  1. 1.
    $h_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],[[\omega_{k}^{\prime}]])=\frac{1}{(2i\pi)^{r}}(2\pi)^{\nu_{k_{0}}+r}a_{k,\overline{k}}c_{k,\mu+n-k}\frac{1}{q^{q}}S^{o}(\omega_{0},\omega_{n})$ if $k\neq 0$,
  2. 2.
    $h_{\mathcal{D},n}([[\omega_{0}^{\prime}]],[[\omega_{0}^{\prime}]])=\frac{1}{(2i\pi)^{n}}(2\pi)^{2n}a_{0,0}S^{o}(\omega_{0},\omega_{n})$,
  3. 3.
    $h_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],[[\omega_{\ell}^{\prime}]])=0$ for any primitive element $\omega_{\ell}^{\prime}]$ of weight $\nu_{k_{0}}$ such that $\ell\neq k$.
Proof.  By Proposition B.2, we have $h_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],[[\omega_{k}^{\prime}]])=i^{2p-(r+\nu_{k_{0}})}g_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],\overline{[[\omega_{k}^{\prime}]]})$ where $p=r-[\alpha_{k}]$ hence, using (53),
$h_{\mathcal{D},\nu_{k_{0}}}([[\omega_{k}^{\prime}]],[[\omega_{k}^{\prime}]])=i^{2r-2[\alpha_{k}]-r}(-1)^{r-[\alpha_{k}]}(2\pi)^{\nu_{k_{0}}}a_{k,\overline{k}}c_{k,\mu+n-k}\frac{1}{q^{q}}S^{o}(\omega_{0},\omega_{n})$
$i^{-r}(2\pi)^{\nu_{k_{0}}}a_{k,\overline{k}}c_{k,\mu+n-k}\frac{1}{q^{q}}S^{o}(\omega_{0},\omega_{n}).$
This shows the first point, and the computations are analogous for the second one. The third point also follows from Proposition B.2.   $\Box$
Remark B.4
Proposition B.3 shows that it is natural to consider the bilinear form $(2i\pi)^{r}g_{\mathcal{D}}$. The bilinear form $(2i\pi)^{-r}g_{\mathcal{D}}$ is basically the orbifold Poincaré-Saito duality defined in Section 5.1 (the Poincaré-Lefschetz conjugation (24) corresponds to the case $a_{k,\overline{k}}=1$ in Lemma B.1): this shift from $(2i\pi)^{r}g_{\mathcal{D}}$ to $(2i\pi)^{-r}g_{\mathcal{D}}$ is due to the fact that we replace $(-1)^{r-[\alpha_{k}]}$ by $(-1)^{[\alpha_{k}]}$ in Formula (53).